### v1

# DEFT: Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time

Abstract
Autonomous wire harness assembly requires robots to manipulate complex branched cables with high precision and reliability. A key challenge in automating this process is predicting how these flexible and branched structures behave under manipulation. Refer to caption Figure 1: This paper introduces DEFT, a novel framework that combines a differentiable physics-based model with a learning-based approach to accurately model and predict the dynamic behavior of Branched Deformable Linear Objects (BDLOs) in real time. As this paper illustrates, this model can be used in concert with a motion planning algorithm to autonomously manipulate BDLOs. The figures above illustrate how DEFT can be used to autonomously complete a wire insertion task. Top: The system first plans a shape-matching motion, transitioning the BDLO from its initial configuration to the target shape (contoured with yellow), which serves as an intermediate waypoint. Bottom: Starting from the intermediate configuration, the system performs thread insertion, guiding the BDLO into the target hole while also matching the target shape. Notably, DEFT predicts the state of the wire recursively without relying on ground truth or perception data at any point in the process. Without accurate predictions, it is difficult for robots to reliably plan or execute assembly operations. While existing research has made progress in modeling single-threaded Deformable Linear Objects (DLOs), extending these approaches to Branched Deformable Linear Objects (BDLOs) presents fundamental challenges. The junction points in BDLOs create complex force interactions and strain propagation patterns that cannot be adequately captured by simply connecting multiple single-DLO models. To address these challenges, this paper presents Differentiable discrete branched Elastic rods for modeling Furcated DLOs in real-Time (DEFT), a novel framework that combines a differentiable physics-based model with a learning framework to: 1) accurately model BDLO dynamics, including dynamic propagation at junction points and grasping in the middle of a BDLO, 2) achieve efficient computation for real-time inference, and 3) enable planning to demonstrate dexterous BDLO manipulation. A comprehensive series of real-world experiments demonstrates DEFT’s efficacy in terms of accuracy, computational speed, and generalizability compared to state-of-the-art alternatives.
I Introduction
The automation of wire harness assembly represents a critical challenge in manufacturing, particularly in the automotive and aerospace industries where complex, branched cable systems are essential components. These assemblies require robots to manipulate Branched Deformable Linear Objects (BDLOs) with precision over extended periods ($\geq$5s) [1, 2, 3], a task that remains largely unautomated due to the objects’ complex nonlinear dynamics. While accurate prediction of BDLO behavior is possible through computationally intensive models, real-time manipulation demands rapid computation. Further complicating matters, practical assembly environments often involve occlusions that prevent perception systems from reliably estimating the full BDLO configuration.
Existing approaches to modeling BDLOs  [4, 5, 6, 7] have focused primarily on computer graphics applications, prioritizing visual plausibility over physical accuracy. These models typically require extensive manual parameter tuning and make impractical assumptions about perfect state observation—limitations that render them unsuitable for robotic manipulation tasks. Recent advances in modeling single Deformable Linear Objects (DLOs) [8, 9, 10, 11, 12] have shown promise, but the extension to branched structures introduces fundamental new challenges. When a robot manipulates one branch of a wire harness, the forces and deformations propagate through junction points to affect all connected branches. This coupling between branches, combined with the need for flexible grasp locations and real-time computation, makes BDLO manipulation significantly more complex than handling single DLOs.
To address these challenges, this paper introduces Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time (DEFT). This framework introduces a novel junction coupling approach that accurately models dynamic propagation between parent and child branches while maintaining real-time performance. As this paper illustrates through real-world evaluations, DEFT enables three essential wire harness assembly tasks: 3D shape matching, multi-thread grasp, and precise thread grasping and insertion. The contributions of this paper are four-fold:
  1. 1.
    A novel differentiable BDLO model that accurately captures how dynamics propagate through wire harness junction points thereby enabling model parameters to be efficiently identified via experimental observation;
  2. 2.
    A computational representation that enables efficient inference and model parameter learning through parallel programming in concert with analytical gradients;
  3. 3.
    A specialized Graph Neural Network (GNN) architecture for BDLO residual learning, which can compensate for numerical integration errors that arise over long time horizon prediction;
  4. 4.
    A comprehensive experimental validation demonstrating DEFT’s superior performance in accuracy, speed, and sampling efficiency compared to existing approaches, including integration into a motion planning framework for real-world manipulation tasks.
The framework is implemented in PyTorch for seamless integration with deep learning workflows, and represents, to the best of our knowledge, the first quantitative investigation of BDLO modeling and manipulation in real-world settings. The results demonstrate that DEFT enables reliable autonomous manipulation of complex wire harnesses, addressing a significant challenge in manufacturing automation. The code and dataset associated with this paper can be found at https://github.com/roahmlab/DEFT.
II Related Work
Effective BDLO modeling requires three critical capabilities: real-time prediction, accurate 3D simulation, and efficient model parameter learning. Real-time prediction capability is essential for robotic manipulation tasks, where quick decision-making can mean the difference between successful and failed object handling. Three-dimensional accuracy ensures the model can capture the complex spatial deformations that BDLOs undergo during manipulation, including twisting, bending, and branch interactions. Efficient parameter learning enables models to automatically adapt to different materials and configurations through data-driven optimization, reducing both computational overhead and the need for extensive training data. This section reviews relevant approaches in modeling BDLOs and single DLOs while considering each of these criteria.
II-A DLOs Modeling
Approaches to modeling DLOs have evolved from purely physics-based methods toward more sophisticated learning-based frameworks that balance between the three aforemention key capabilities.
Physics-based approaches implement different mathematical frameworks to model physical behavior. Mass-spring systems represent objects as networks of point masses connected by springs, computing forces and accelerations to simulate movement [13, 14, 15]. Position-Based Dynamics (PBD) directly manipulates vertex positions to satisfy physical constraints, avoiding expensive force calculations [10, 16, 9]. Finite Element Methods (FEM) divide objects into discrete elements and solve partial differential equations to compute deformations [17, 18, 19]. Discrete Elastic Rods (DER) models represent objects as centerline curves with material frames, enabling explicit modeling of bending and twisting [16, 20, 21, 22]. Unfortunately each of these methods has limitations: mass-spring systems describe the rigidity of wires by introducing large spring stiffness which can make prediction numerically unstable; PBD models struggle with parameter learning because they assume that the system is quasi-static; FEM methods are computationally intensive; and DER models lack differentiable expressions for their model parameters, making efficient parameter optimization challenging.
Learning-based modeling approaches use data-driven techniques to predict deformation behavior. Methods such as Graph Neural Networks (GNN) process vertex-level spatial information [12], while Bi-LSTM models capture temporal dependencies in deformation sequences [11]. These approaches excel at real-time prediction through efficient computation and support straightforward parameter learning through standard optimization techniques. However, they struggle with 3D accuracy, particularly in capturing the complex dynamic deformations characteristic of real-world manipulation.
Recent work has explored hybrid approaches that combine learning with physics-based modeling. This is exemplified by DEFORM [8] which constructs a novel differentiable DER model and combines it with a neural network-based residual learning approach to account for numerical integration errors during prediction. This hybrid approach achieves both real-time performance and 3D accuracy while enabling efficient parameter learning through end-to-end differentiability.
Despite DEFORM’s advances in DLO modeling, it faces three significant limitations when extended to BDLO applications. First, DEFORM’s modeling framework does not address the dynamics of branched structures, particularly the propagation of stretching, twisting, and bending forces at junction points. Second, DEFORM’s manipulation framework assumes fixed clamping points at the object ends, following conventional DLO literature. This assumption restricts its applicability to more sophisticated manipulation tasks, such as thread insertion, which require grasping at multiple points along the object’s length. Third, DEFORM’s computational approach presents efficiency challenges for BDLO simulation. Its sequential processing of individual branches would result in linear computational time increases as branches are added. Furthermore, its bi-level optimization framework, while enabling learning from real-world data, relies on numerical gradients that can lead to slow convergence and optimization instability.
II-B BDLOs Modeling
Existing BDLO modeling approaches adapt various physical simulation frameworks to handle branched structures. FEM-based methods extend traditional finite element analysis by treating branches as connected elastic elements [23]. While these methods excel in 3D accuracy and enable efficient parameter learning through gradient-based optimization, their computational requirements preclude real-time prediction. DER-based approaches introduce rigid-body constraints at junctions to ensure proper alignment of attached segments  [4], achieving real-time performance and reasonable 3D accuracy but struggling with parameter learning due to their fixed parameter structure. Quasi-static methods like ASMC enforce geometric constraints while assuming static equilibrium conditions [5], offering real-time performance, but sacrificing 3D accuracy by overlooking dynamic effects that can make parameter learning challenging.
To address these challenges, DEFT introduces a novel differentiable framework specifically designed for BDLOs. Through a specialized GNN architecture for branched structures, DEFT achieves all three critical capabilities: real-time performance, 3D accuracy, and efficient parameter learning. This enables DEFT to adapt to various materials and configurations while maintaining the speed and accuracy necessary for practical robotic manipulation tasks.
III Preliminaries
This section introduces the key components involved in modeling DLOs. It then summarizes the primary contributions of DEFORM and describes the challenges of adapting it to BDLO modeling.
III-A Deformable Linear Object (DLO) Model
To model a DLO, we represent it as an indexed set of $n$ vertices at each time instance. The ground truth location of vertex $i$ at time $t$ is denoted by $\mathbf{x}_{t}^{i}\in\mathbb{R}^{3}$, and the set of all $n$ vertices is denoted by $\mathbf{X}_{t}$. A segment, or edge, in the DLO is the line between successive vertices, $\mathbf{e}^{i}_{t}=\mathbf{x}^{i+1}_{t}-\mathbf{x}^{i}_{t}$. Let $\mathbf{E}_{\text{t}}$ correspond to the set of all edges at time $t$. Let $\mathbf{M}_{i}\in\mathbb{R}^{3\times 3}$ denote the mass matrix of vertex $i$. The velocity of the vertices of the DLO is approximated by $\mathbf{V}_{t}=(\mathbf{X}_{t}-\mathbf{X}_{t-1})/{\Delta t}$, where $\Delta t$ is the time step between frames. Note that $\mathbf{V}_{t}$ is an approximation of the actual velocities. We distinguish between ground truth elements and predicted elements by using the circumflex symbol (e.g., $\mathbf{X}_{t}$ is the ground truth set of vertices at time $t$, and $\hat{\mathbf{X}}_{t}$ is the predicted set of vertices at time $t$).
III-B Modeling DLOs with DEFORM
III-B1 Discrete Elatic Rods (DER)
DEFORM builds upon DER theory to model three distinct deformation modes of DLOs: bending, twisting, and stretching. DER theory models DLOs using two key families of coordinate frames. First, it introduces Bishop Frames, which provide a twist-free reference state along the centerline of the rod. The Bishop Frame only rotates to follow the curve’s geometry, providing a relaxed baseline configuration for the DLO. Second, the theory employs Material Frames, which describe the actual physical deformation of the rod. Material Frames can be expressed relative to the Bishop Frames through a single angle whose rate of change represents the physical twist of the DLO. To describe the twist of the DLO, DER theory assumes that the angle of the Material Frame with respect to the Bishop Frame is one that minimizes the potential energy of the DLO. We briefly summarize DER theory in this section, but a longer introduction can be found here [24].
Bishop Frame: The Bishop Frame at edge $i$ is made up of three axes $\{\mathbf{t}^{i},\mathbf{b}_{1}^{i},\mathbf{b}_{2}^{i}\}$. We next formally define each axis of the Bishop Frame. First, $\mathbf{t}^{i}\in\mathbb{R}^{3}$ is the unit tangent along edge $i$ which is defined as $\mathbf{t}^{i}=(\mathbf{x}^{i+1}-\mathbf{x}^{i})/||\mathbf{x}^{i+1}-\mathbf{x}^{i}||_{2}$. The vector $\mathbf{b}_{1}^{i}\in\mathbb{R}^{3}$ is defined iteratively by applying a rotation matrix to the previous Bishop frame: $\mathbf{b}_{1}^{i}=\mathbf{R}^{i}\cdot\mathbf{b}_{1}^{i-1}$. The rotation matrix $\mathbf{R}^{i}\in\mathbb{R}^{3\times 3}$ satisfies the following conditions
$\mathbf{t}^{i}=\mathbf{R}^{i}\cdot\mathbf{t}^{i-1}$ (1)
$\mathbf{t}^{i-1}\times\mathbf{t}^{i}=\mathbf{R}^{i}\cdot(\mathbf{t}^{i-1}\times\mathbf{t}^{i}).$ (2)
The first condition ensures that the tangent vector of the DLO at each subsequent edge can be described using the rotation matrix and the tangent vector at the previous edge. Though we do not prove it here, the second condition ensures that the rotation between successive Bishop Frames is minimal while moving along the DLO. In particular, this latter condition ensures that there is no twist about the subsequent tangent vectors at each edge of the DLO. Finally, $\mathbf{b}_{2}^{i}:=\mathbf{t}^{i}\times\mathbf{b}_{1}^{i}$, which ensures that $\mathbf{b}_{2}^{i}$ is orthogonal to $\mathbf{t}^{i}$ and $\mathbf{b}_{1}^{i}$.
Material Frame: Subsequently, DER theory introduces Material Frames that are made up of three axes $\{\mathbf{t}^{i},\mathbf{m}_{1}^{i},\mathbf{m}_{2}^{i}\}$. These Material Frames are generated by rotating the Bishop frame about the tangent vector by a scalar $\theta^{i}$, i.e., $\mathbf{m}_{1}^{i},\mathbf{m}_{2}^{i}$ are defined as
$\mathbf{m}_{1}^{i}=\textbf{b}_{1}^{i}\cos\theta^{i}+\textbf{b}_{2}^{i}\sin\theta^{i}$ (3)
$\quad\mathbf{m}_{2}^{i}=-\textbf{b}_{1}^{i}\sin\theta^{i}+\textbf{b}_{2}^{i}\cos\theta^{i}$ (4)
An illustration of the frames can be found in Figure 2.
Figure 2: An illustration of DER coordinate frames.
Equations of Motion: Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ denote the set of all $n-1$ Material Frames at time $t$. Let $\bm{\alpha}$ represent the vector of material properties for each vertex of the DLO (i.e., mass, bending modulus, and twisting modulus) [4, Section 4.2]. Using these definitions, one can compute the potential energy of the DLO arising from bending and twisting which we denote by $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$. DER theory assumes that the DLO reaches an equilibrium state between each simulation time step to obtain the optimal $\bm{\theta}_{t}^{*}$:
$\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ (5)
Once $\bm{\theta}^{*}_{t}(\mathbf{X}_{t},\bm{\alpha})$ is derived, the restorative force during deformation is given by the negative gradient of the potential energy with respect to the vertices. Consequently, the equation of motion for the DLO is:
$\mathbf{M}\ddot{\textbf{X}}_{t}=-\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}}$ (6)
One can numerically integrate this formula to predict the velocity and position by applying the Semi-Implicit Euler method:
$\hat{\textbf{V}}_{t+1}$ $=\hat{\textbf{V}}_{\text{t}}-\Delta_{\text{t}}\mathbf{M}^{-1}\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}},$ (7)
$\hat{\textbf{X}}_{\text{t}+1}$ $=\hat{\textbf{X}}_{\text{t}}+\Delta_{\text{t}}\hat{\textbf{V}}_{t+1},$ (8)
where $\Delta_{\text{t}}>0$ is a user specified time discretization.
III-B2 Differentiable Discrete Elastic Rods (DDER)
Tuning the DER model parameters, $\bm{\alpha}$, to match real-world DLOs is non-trivial. To address this challenge, DEFORM introduces DDER, which enables the automatic tuning of DLO parameters using real-world data. Predicting $\hat{\textbf{X}}_{t+1}$ with DDER through (7) and (8) involves solving an optimization problem as defined in (5). Consequently, minimizing the prediction loss requires solving a bi-level optimization problem at each time step $t$:
$\underset{\bm{\alpha}}{\min}\hskip 14.22636pt$ $\|\textbf{X}_{t+1}-\hat{\textbf{X}}_{t+1}(\bm{\theta}^{*}_{t})\|_{2}\hskip 42.67912pt$ (9)
$\bm{\theta}^{*}_{t}=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ (10)
Note $\hat{\textbf{X}}_{t+1}$ depends on the previously computed $\bm{\theta}^{*}_{t}$ through (6), (7) and (8). To compute the gradient of (10) with respect to $\bm{\alpha}$, DEFORM employs solvers from Theseus [25], which use PyTorch’s automatic differentiation [26]. As we illustrate via experiments in Section V, this can be a time-consuming process.
III-B3 Integration Method with Residual Learning
To compensate for numerical errors that arise during numerical integration(8), DEFORM incorporates a learning-based framework into the integration method:
$\hat{\textbf{X}}_{t+1}=\hat{\textbf{X}}_{t}+\Delta_{t}(\hat{\textbf{V}}_{t+1}+\text{GNN}(\hat{\textbf{X}}_{t},\hat{\textbf{V}}_{t},\bm{\alpha}))$ (11)
This structure builds on residual learning [27], grounding the neural network in physical laws while leveraging data-driven insights.
III-B4 Inextensiblity Enforcement with Momentum Conservation
After numerical integration, the DLO length may deviate from its initial value due to numerical errors, making it difficult to accurately model inextensible DLOs. A common physics-based approach [13, 14, 15] uses mass-spring systems to minimize this deviation, but this can result in stiff simulations and numerical instability. DEFORM addresses this by enforcing an inextensibility constraint while preserving momentum, leading to more accurate and stable simulation.
III-B5 Challenges of Extending DEFORM to BDLOs
Though DEFORM can accurately simulate single-thread DLOs, it is unsuitable for BDLOs for several reasons. First, the definitions of Bishop and Material Frames do not generalize to BDLOs. To see why this is true, recall that at junctions in a BDLO there are several tangent vectors because there are more than two edges. As a result, satisfying (1) and (2) results in an overdetermined system which makes it impossible to find a rotation matrix $\mathbf{R}^{i}$ that satisfies all constraints. Consequently, the Bishop and Material Frames cannot be defined at these junctions. An alternative to extend DEFORM to BDLOs is to simulate each branch individually using DEFORM. However, this requires repeatedly solving (5), which reduces computational efficiency. This is further complicated when trying to enforce inextensibility using the momentum conservation constraint. To address this, DEFT proposes a computational representation that supports parallel processing, enabling more efficient inference.
Second, relying on automatic differentiation to perform parameter tuning becomes computationally prohibitive when dealing with BDLOs. To address this limitation, DEFT utilizes the analytical gradients of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$. Third, the GNN used by DEFORM during residual learning is tailored to single-strand DLOs. To address this limitation, DEFT introduces a specialized GNN Architecture for residual learning for BDLOs that can compensate for numerical integration errors over long horizon predictions.
Figure 3: Algorithm Overview of DEFT. In the initialization stage (Section IV-A), DEFT begins by separating the BDLO into a parent DLO and one or more children DLOs. Each DLO is discretized into vertices and represented as elastic rods. This setup allows DEFT to capture the geometric and physical properties required for dynamic simulation. To improve computational efficiency, DEFT then predicts the dynamics of each branch in parallel (Section IV-F). During this process, analytical gradients are provided to solve (10), as detailed in Section IV-C, ensuring efficient and stable convergence. Next, to address numerical errors, DEFT employs a GNN designed to learn the BDLO’s residual dynamics (Section IV-D). By modeling discrepancies between simulated and observed behavior, the GNN refines predictions and enhances overall accuracy. After integration, DEFT enforces constraints (Section IV-E) to enforce physical realism. Inextensibility constraints are applied to each branch, while junction-level constraints ensure proper attachment at branch junctions. Additionally, edge orientation constraints enable the propagation of dynamics across these junctions. Throughout the entire pipeline, all components remain fully differentiable, allowing for efficient parameter learning from real-world data.
IV Methodology
Our objective is to model the dynamic behavior of a BDLO under robotic control. Traditional modeling approaches simplify the problem by assuming that the DLO’s endpoints are each held by separate robot end-effectors [8, 28, 16, 10, 29, 30]. Although this assumption reduces complexity, it limits the potential for dexterous manipulation (e.g., inserting the middle of a BDLO into a plug). This work relaxes this assumption and considers scenarios where the BDLO is grasped at an intermediate point, rather than exclusively at its ends.
This section introduces the key components of BDLO modeling and overviews the various components of DEFT, which are illustrated in Figure 3. The key components of DEFT are: 1) analytical gradients to accelerate the optimization procedure used to solve (5); 2) a GNN architecture tailored for residual learning in BDLO simulations; 3) an inextensible constraint formulation that enables the propagation of dynamics at the junction between branches while conserving momentum, which improves the numerical stability of the simulation; and 4) a parallel programming strategy that ensures efficient simulation and gradient computation during training and simulation. Please note that to simplify the presentation, we omit the time subscript $t$ in Sections IV-C and IV-E.
IV-A Branched Deformable Linear Object Model
Figure 4: An illustration of a branched deformable linear object (BDLO) model. Left: An example configuration of BDLO manipulation. Right: Relevant notation and structure at a junction.
To model a BDLO, we decompose its structure into a hierarchical arrangement of DLOs. The parent DLO is defined as the minimum-length path through the BDLO that connects the two end-effectors (i.e., the points where the structure is manipulated or anchored). All other branches diverging from this primary path are modeled as children DLOs that maintain their physical connection to the parent DLO at their attachment points, which we refer to as junction points. The parent and children DLOs are each modeled as DLOs and inherit all previously introduced notation. For conciseness, we describe the junction point using only three edges at each time instance: two from the parent DLO and one from a child DLO. In principle, the methods in DEFT can be extended to handle junctions with more than three edges.
We distinguish between the parent and child DLOs using superscripts $p$ and $c$, respectively. For instance, if a junction is located at the $i$-th vertex, the associated edges are denoted by $\mathbf{e}^{p,i-1}_{t}$, $\mathbf{e}^{p,i}_{t}$, and $\mathbf{e}^{c,1}_{t}$. An illustration of a BDLO can be found in Figure 4. To capture the influence of bending and twisting, we introduce the moment of inertia of the $i$-th edge which is denoted by $\mathbf{I}^{\text{i}}\in\mathbb{R}^{3\times 3}$. It represents how the segment’s mass distribution affects its resistance to rotational deformations. The orientation of $\mathbf{e}^{i}_{t}$ is denoted by an angle axis vector $\mathbf{\Omega}^{i}_{t}\in\mathbb{R}^{3}$.
Let the orientation and position of the two end effectors at time $t$ be denoted by $\mathbf{u}_{t}\in\mathbb{R}^{12}$. We refer to $\mathbf{u}_{t}$ as the input. We assume that the initial state of the BDLO is fully observable, and that it can be accurately estimated using image segmentation [31] in combination with Gaussian Mixture Models [32].
IV-B DEFT Overview
At each time $t$, DEFT simulates the behavior of a BDLO when given predictions of the vertices and their velocities under a specified input as described in Algorithm 1. It begins by first associating the grasped vertices with the robot end effector inputs $\mathbf{u}_{t}$. DEFT then processes the parent and children DLOs in parallel to compute their respective material frames using (3) (4), which are subsequently used to calculate each branch’s potential energy. DEFT then minimizes this potential energy using analytical gradients to determine the optimal parameters $\bm{\theta}_{t}^{*}$ and updates the potential energy accordingly (Line 2). These forces are integrated forward in time using a Semi-Implicit Euler method combined with residual learning to update the vertex velocities and positions (Lines 3 and 4).
However, due to numerical integration errors, branch lengths may change in length. In addition, up to this point, DEFT treats each branch’s internal dynamics in isolation, without accounting for junction interactions. This omission can hinder the propagation of dynamics across junctions. For example, children DLOs may detach from their parent branches, or deformation effects (e.g., bending and twisting) may not propagate through the junction. To address these issues, DEFT enforces additional constraints to ensure each branch remains inextensible, maintain branch connections, and allow deformation effects to propagate across the junction (Line 5). By applying Algorithm 1 over multiple time increments, DEFT can accurately predict the BDLO’s state over extended time horizons. An algorithm overview is illustrated in Figure 3.
Algorithm 1 $(\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}})=\textbf{DEFT}(\hat{\mathbf{X}}_{\text{t}},\hat{\mathbf{V}}_{\text{t}},\textbf{u}_{t}$)
1:Associate grasped vertices with $\textbf{u}_{t}$
2:Calculate Material Frames $\Gamma(\textbf{X}_{t},\bm{\theta}_{t})$triangleright (3),(4)
3:$\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$triangleright Section.IV-C
4:$\hat{\textbf{V}}_{t+1}=\hat{\textbf{V}}_{t}-\Delta_{t}\mathbf{M}^{-1}\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}}$triangleright (7)
5:$\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t}+\Delta_{t}\hat{\textbf{V}}_{t+1}+\text{GNN}$triangleright Section.IV-D
6:Constraints enforcement on $\hat{\mathbf{X}}_{t+1}$triangleright Algorithm 2
7:$\hat{\textbf{V}}_{t+1}=(\hat{\mathbf{X}}_{t+1}-\hat{\mathbf{X}}_{t})/\Delta_{t}$triangleright Velocity Update
return $\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}}$
IV-C Analytical Gradient Derivation
The goal of this subsection is to present the analytical gradient of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$. To improve the paper’s readability, we summarize the result in the following theorem whose proof can be found in Appendix -A.
Theorem 1.
Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ be the Material Frame at time $t$, and let $\bm{\alpha}$ contain the bending stiffness $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and twisting stiffness $\beta\in\mathbb{R}$. Suppose the total potential energy is the sum of bending and twisting potential energies:
$\begin{split}P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})\;=\;P_{\text{bend}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\mathbf{B}\bigr{)}+\\ P_{\text{twist}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\beta\bigr{)}.\end{split}$ (12)
Then the partial derivative of the bending and twisting potential energies with respect to $\theta^{i}$ are given by
$\begin{split}\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(&\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\\ &\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},\end{split}$ (13)
$\begin{split}\frac{\partial P_{\mathrm{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\beta^{i}\,\bigl{(}\theta^{i}-\theta^{i-1}\bigr{)}\;-\;\\ \beta^{i+1}\,\bigl{(}\theta^{i+1}-\theta^{i}\bigr{)},\end{split}$ (14)
respectively, where the material curvature $\bm{\omega}^{(i,j)}$ is defined as
$\bm{\omega}^{(i,j)}\;=\;\bigl{(}\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{1}^{j},\;\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{2}^{j}\bigr{)}^{T}\quad\text{for}\quad j\in\{i-1,\,i\},$ (15)
$\overline{\bm{\omega}}^{(i,j)}$ denotes the undeformed material curvature, and the curvature vector $\kappa\mathbf{b}^{i}$ in (15) is
$\kappa\,\mathbf{b}^{i}\;=\;\frac{2\,\bigl{(}\mathbf{e}^{i-1}\times\mathbf{e}^{i}\bigr{)}}{\|\mathbf{e}^{i-1}\|_{2}\,\|\mathbf{e}^{i}\|_{2}\;+\;\mathbf{e}^{i-1}\cdot\mathbf{e}^{i}}.$ (16)
IV-D Integration Method with Residual Learning
This subsection constructs the GNN for residual learning in (11). The motivation for employing a GNN lies in its ability to represent a BDLO as a graph, where each vertex is treated as a node and edges capture both intra-branch connections and junctions. This graph-based formulation translates into an adjacency matrix $\mathbf{A}\in\mathbb{R}^{(n^{p}+n^{c})\times(n^{p}+n^{c})}$, which links individual branch nodes and explicitly encodes node information across branch junctions. This adjacency matrix is defined as:
$\mathbf{A}_{i_{1},i_{2}}=\begin{cases}1&\text{if there is an edge between }i_{1}\text{ and }i_{2},\\ 0&\text{otherwise,}\end{cases}$ (17)
An illustration of $\mathbf{A}$ can be found in Figure 5. In practice, DEFT constructs $\mathbf{A}$ for each BDLO whose behavior a user wants to predict. This adjacency matrix is utilized within a Graph Convolutional Network (GCN) [33] to aggregate learned features. A single GCN layer is defined as
$\text{GCN}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})=\mathbf{\tilde{A}}\,\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\,\mathbf{W},$ (18)
where $\mathbf{\tilde{A}}$ is the normalized adjacency matrix
$\mathbf{\tilde{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (19)
Note that $\mathbf{\tilde{A}}$ is the normalization of $\mathbf{A}$ using the degree matrix $\mathbf{D}=\mathrm{diag}(d_{1},d_{2},\ldots,d_{n})$ where $d_{i1}=\sum_{i2}A_{i1,i2}$. This normalization ensures that nodes with a high degree do not disproprtionately influence the overall performance. This ensures consistent feature magnitudes across the learned representation [34]. The node feature matrix, $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\in\mathbb{R}^{(n^{p}+n^{c})\times h}$, is constructed by concatenating its arguments together. The trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ is implemented via a multi-layer perceptron (MLP) [35], which projects input features into latent representations.
Figure 5: The adjacency matrix $\mathbf{A}$ for the junction in Figure 4 captures both self-loops (diagonal entries) and inter-node connections (off-diagonal entries). By embedding each node’s local dynamics alongside its coupling to neighboring nodes, $\mathbf{A}$ enables an graph representation of the BDLO’s behavior.
IV-E Enforcing Constraints
After performing numerical integration, DEFT applies constraints to ensure branch inextensibility, ensure that branches are attached at the junctions, and ensure that deformations are propagated throughout the BDLO. To accomplish this, DEFT updates the positions of each vertex using correction terms $\Delta\hat{\mathbf{X}}_{\text{}}$. One could just update the positions of each vertex by just projecting directly onto the constraint set. However, this induces highly nonlinear changes in momentum and can lead to simulation instability and unrealistic behavior.
Motivated by the PDB method [36], DEFT enforces the aforementioned constraints while trying to preserve momentum. DEFT does this by formulating the computation of the correction term as the solution to an optimization problem. As we illustrate through our experimental evaluation in Section V, this leads to more stable simulations, which corresponds to more realistic predictions. Finally, note that DEFT could formulate the optimization problem to compute the correction terms as a single large optimization problem that considers constraints between all vertices in the BDLO simultaneously. However, this large optimization problem can be difficult to solve rapidly. Instead DEFT enforces these constraints sequentially between neighboring nodes. Next, we summarize each of these optimization problems.
IV-E1 Inextensibility Enforcement
To enforce inextensibility of each branch, we define a constraint function between successive vertices. This function penalizes deviations of each segment’s length from its undeformed length. By applying the following theorem, whose proof can be found in [8, Appendix B.2], one can compute $\Delta\hat{\mathbf{x}}^{i},\Delta\hat{\mathbf{x}}^{i+1}$ explicitly:
Theorem 2 (Computing the Corrections to Enforce Inextensibility).
Consider two successive vertices $i$ and $i+1$ along the same branch. Let $\overline{\mathbf{e}}_{i}$ denote the edge between these vertices when the BDLO is undeformed. Let the constraint function to enforce inextensibility be defined as:
$C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=||(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})-(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})||_{2}-||\overline{\textbf{e}}_{i}||_{2}.$ (20)
The correction $\Delta\hat{\mathbf{X}}$ to enforce inextensibility at joint $i$ can be computed by solving the following optimization problem:
$\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (21)
s.t. $\hskip 8.5359pt\Delta\hat{\mathbf{x}}^{j}=0$ $\forall j\notin\{i,i+1\},$
$\hskip 8.5359pt\mathbf{M}^{i}\Delta\hat{\mathbf{x}}^{i}+\mathbf{M}^{i+1}\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{0},$
$\hskip 8.5359pt\mathbf{I}^{i}\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0},$
where $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\in\mathbb{R}^{3}$ is defined in Appendix -B and represents the orientation change of the $i$th edge arising from $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ and $\mathbf{0}$ correspond to the zero vector of appropriate size. The only non-zero elements of the minimizer to this optimization problem can be computed explicitly and are equal to:
$\Delta\hat{\mathbf{x}}^{i}=\mathbf{M}^{i+1}\bigl{(}\mathbf{M}^{i}+\mathbf{M}^{i+1}\bigr{)}^{-1}C^{i}_{I}(\hat{\mathbf{X}},\mathbf{0})\frac{\bigl{(}\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}\bigr{)}}{\|\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}\|_{2}},$ (22)
$\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{M}^{i}(\mathbf{M}^{i}+\mathbf{M}^{i+1}\bigr{)}^{-1}C^{i}_{I}(\hat{\mathbf{X}},\mathbf{0})\frac{\bigl{(}\hat{\mathbf{x}}^{i}-\hat{\mathbf{x}}^{i+1}\bigr{)}}{\|\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}\|_{2}}.$ (23)
Note that the inextensibility constraint at vertex $i$ is satisfied when $C^{i}_{I}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$. For convenience, let the function that computes the minimizer of optimization problem (21) using $\hat{\mathbf{X}}$ be denoted by $D^{i}_{I}$ (i.e., $D^{i}_{I}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
IV-E2 Enforcing Attachment at the Junction
To ensure the child branch remains attached to the parent branch, we define a constraint function between theirjunction vertices. By applying the following theorem, whose proof can be found in Appendix -C, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
Theorem 3.
Consider the parent-branch vertex $\hat{\mathbf{x}}^{p,i}$ and child-branch vertex $\hat{\mathbf{x}}^{c,1}$ at the junction such as the one depicted in Figure 4. Let the constraint function to enforce attachment at the junction be defined as:
$C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=||\bigl{(}\hat{\mathbf{x}}^{p,i}+\Delta\hat{\mathbf{x}}^{p,i}\bigr{)}-\bigl{(}\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1}\bigr{)}||_{2}.$ (24)
Suppose that the correction $\Delta\hat{\mathbf{X}}$ to account for the junction attachment constraint is obtained by solving an optimization problem that is identical to (21) with the objective function replaced with (24) and with a requirement that the only elements of $\Delta\hat{\mathbf{X}}$ are those associated vertices $p,i$ and $c,1$. The only non-zero elements of the minimzer to this optimization problem can be computed explicitly and are equal to:
$\Delta\hat{\mathbf{x}}^{p,i}=\mathbf{M}^{c,1}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}},$
$\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{M}^{p,i}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}\ C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{p,i}-\hat{\mathbf{x}}^{c,1})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}}.$
Note that the junction attachment constraint is satisfied at vertex $p,i$ when $C^{p,i}_{I}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$. For convenience, let the function that computes the minimizer of optimization problem (21) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{A}$ (i.e., $D^{p,i}_{A}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
IV-E3 Edge Orientation Constraints at the Junction
Many practical wire-harnesses feature junctions that are nearly rigid. This near-rigidity arises because manufacturers often use rigid plastics or other stiff materials at these connection points [37]. This results in minimal rotational or bending freedom. Consequently, the dynamics between the parent and child branches propagate through the junction in an almost rigid manner. To capture this behavior, we model each edge as a rigid body and define a constraint function that computes the relative orientation of each body. We then require that this constraint function is approximately satisfied where ap By applying the following theorem, whose proof can be found in Appendix -D, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
Theorem 4.
Consider the parent-branch vertex $\hat{\mathbf{x}}^{p,i}$ and child-branch vertex $\hat{\mathbf{x}}^{c,1}$ at the junction such as the one depicted in Figure 4. Let $\epsilon>0$ be a user specified parameter. Let the constraint function to enforce the orientation of the junction be defined as:
$C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\big{)}-\\ \big{(}\hat{\bm{\Omega}}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}}\big{)}\|_{2}-\epsilon,$ (25)
where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix -B and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid. The orientation correction at the junction is obtained by solving the following optimization problem:
$\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (26)
s.t. $\hskip 8.5359pt\mathbf{I}^{p,i}\hat{\bm{\Omega}}_{\Delta}^{p,i}+\mathbf{I}^{c,1}\hat{\bm{\Omega}}_{\Delta}^{c,1}=\mathbf{0}$
When $\Delta_{t}$ is sufficiently small, then one can compute explicit formulas for $\Delta\hat{\mathbf{X}}$ which can be found in Appendix -D.
Note that the junction orientation constraint is satisfied at vertex $p,i$ when $C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$. For convenience, let the function that computes the minimizer of optimization problem (21) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{O}$ (i.e., $D^{p,i}_{O}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
IV-E4 Algorithm to Implement Corrections Due To Constraint Enforcement
As described earlier, we run each of these optimization problems sequentially to generate a correction to the output of the integration method (Line 5 in Algorithm 1). This is summarized in Algorithm 2. Note that one begins by checking whether any constraint is violated by more than some user-specified parameter $\kappa$. If any of the constraints are violated by more than $\kappa$, then one applies the correction described in the previous subsections to modify the prediction. This repeats until all constraints are satisfied.
Algorithm 2 Momentum Preserving Constraint Enforcement
1:$\hat{\mathbf{X}}_{\text{t+1}},\kappa>0$
2:while $\exists(i,j)\text{ such that }C^{i}_{j}(\hat{\mathbf{X}}_{t+1},\mathbf{0})^{2}>\kappa$ do
3:     for $i=1$ to $n-1$ do
4:         $\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t+1}+D^{i}_{I}(\hat{\mathbf{X}}_{t+1})$triangleright Theorem 2
5:         if $i$th segments is a junction then
6:              $\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t+1}+D^{i}_{A}(\hat{\mathbf{X}}_{t+1})$triangleright Theorem 3
7:              $\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t+1}+D^{i}_{O}(\hat{\mathbf{X}}_{t+1})$triangleright Theorem 4
8:         end if
9:     end for
10:end while
11:return : $\hat{\textbf{X}}_{t}$triangleright Updated Vertices
IV-F Parallel Programming
We implement DEFT using PyTorch [26], taking advantage of its batch operations to execute Algorithm 1. PyTorch’s support for batch-wise computations enables parallel processing of multiple parent and child branch instances, significantly enhancing computational efficiency as we illustrate empirically in Section V-D. Specifically, the first batch is dedicated to executing the parent branch’s algorithm, while the remaining batches handle executions for the children branches.
One challenge arises when the number of nodes in the parent branch ($n^{p}$) differs from the number in a child branch ($n^{c}$). This dimensional mismatch prevents the direct construction of uniform matrices for batch operations. To overcome this, we assume $n^{p}>n^{c}$ and introduce padded matrices and vertices by appending zeros to the child branch’s structure to match the parent branch’s dimensions. This strategy ensures uniform matrix construction across the batches, which facilitates seamless batch-wise operations. During parameter updates, we mask out the gradients from padded elements to prevent them from influencing the learned parameters.
V Experiments and Results
This section presents the real-world evaluation of DEFT, covering: 1) A quantitative accuracy comparison with state-of-the-art methods, including scenarios where grasping occurs at various points along the BDLO, such as parent-branch ends and midpoints; 2) An assessment of DEFT’s inference speed; 3) An ablation study exploring parameter auto-tuning, residual learning, and a naive multi-branch approach that connects multiple single-DLO models; 4) Shape-matching manipulation demonstrations, highlighting DEFT’s real-world manipulation efficacy; 5) Thread insertion manipulation demonstrations, showcasing DEFT’s versatility. Code and data can be found at https://github.com/roahmlab/DEFT. To the best of our knowledge, this will be the first publicly available dataset and code for modeling BDLOs.
V-A Hardware Setup
Figure 6: An illustration of experiment setup. Top: Data collection setup with a dual-arm robot. Bottom: An illustration of the BDLOs used to evaluate and compare the performance of DEFT against baseline methods. To increase dataset diversity, we vary the BDLOs’ stiffness, length, weights, number of vertices, and junction structures.
TABLE I: A summary of each BDLO’s material properties and marker count used in the real-world experiment. Stiffness is estimated based on a relative scale.
BDLO Parent Stiffness / #Markers Child1 Stiffness / #Markers Child2 Stiffness / #Markers Dataset 1 Time (s) Dataset 2 Time (s)
1 Medium / 13 Medium / 4 Medium / 3 505 440
2 Low / 12 Low / 4 Low / 4 735 N/A
3 Medium / 12 Low / 2 High / 3 645 445
4 Medium / 12 High / 2 High / 2 495 N/A
V-A1 Hardware Setup
Four distinct BDLOs were constructed to validate DEFT’s modeling accuracy as shown at the bottom of Figure 6. The material stiffness description of each BDLO is outlined in Table I. Note that we introduced stiff child branches attached to the parent branch in BDLO3 and BDLO4. These branches represent the rigid components (e.g., fixations) commonly used in BDLO construction [37]. These attachments help ensure secure mounting, preserve structural integrity, and manage loads in real‐world applications. Spherical markers were attached to each BDLO to facilitate dataset creation for both training and evaluation, with the number of vertices set equal to the number of MoCap markers. Because the markers are present during both training and evaluation, their impact on the BDLO dynamics is inherently accounted within all subsequent analyses. If one wishes to obtain a BDLO dataset without relying on real-world motion capture, a finite element method-based simulator could be employed. Note that the markers are only used to create training and evaluation datasets, as well as evaluate the performance of DEFT by providing the ground truth. However, for real-world applications, the markers are not needed. For dual manipulation, a Franka Emika Research 3 robot and a Kinova Gen3 robot were used. An illustration of the experimental setup is shown in Figure 6. We used an OptiTrack motion capture system to record ground-truth vertex locations. Note in our implementation of DEFT we set $\epsilon=0.1$ (Theorem 4) and $\kappa=0.02$ (Algorithm 2).
V-A2 Dataset Collection and Training
We collected dynamic trajectory data for selected BDLOs in the real world using a motion capture system operating at 100 Hz, capturing both slow and fast BDLO motions, as illustrated at the top of Figure 6. We split our data collection into two categories:
  • Dataset 1: each robot grasps one end of the BDLO. Data from this category were collected for all BDLOs.
  • Dataset 2: One robot grasped one end of the BDLO while the other robot grasped the BDLO at its midpoint. This dataset is collected for BDLOs 1 and 3.
Total collection times are provided in Table I. The dataset is split with $\sim$75for training and $\sim$25for evaluation. The model for each of the four BDLOs is trained separately using the trajectory data collected from our real-world experiments. During training, models are trained using the same 100 steps (i.e., 1s) of dynamic motion Each method’s performance is evaluated when predicting the state of the wire over a 500 step (i.e. 5s) prediction horizon using just the state of the wire at the initial time and the orientation and position of the two end effectors at each time step.
V-B Baseline Comparisons
Only a limited number of papers have investigated the dynamic behavior of BDLOs, and none provide a differentiable framework for learning real-world dynamics from real-world datasets. Moreover, no public code repositories are available for these approaches, making direct replication infeasible. We select two available implementations, Tree-LSTM [38] and GCN [33] as baseline prediction models for BDLOs. Tree-LSTM and GCN both capture structured relationships. Specifically, Tree-LSTM leverages hierarchical tree representations, while GCN excels in graph-based modeling. Because BDLOs can be described by nodes and edges, these architectures are naturally able to incorporate a BDLO’s dynamics.
V-C Modeling Results
Table II summarizes the average RMSE loss when comparing DEFT to baseline methods for two categories of dataset. Performance is evaluated over a 5 second prediction horizon for four different BDLOs. Note that DEFT outperforms all baselines across all four BDLOs. An illustration of the performance of DEFT, Tree-LSTM, and an ablation of DEFT, which is described in more detail in the next subsection, can be found in Figure 7. A video of related experiments can be found in the supplementary material.
TABLE II: Modeling Results Over a $5$s Horizon
Modeling Accuracy (RMSE, $10^{-2}$ m)
Dataset 1 Dataset 2
Method 1 2 3 4 1 3
GCN [33] 6.70 6.42 5.25 4.20 7.33 4.24
Tree-LSTM [38] 5.32 4.50 2.46 2.71 7.08 4.55
DEFT 1.87 2.82 1.51 1.41 1.49 1.80
Figure 7: Visualization of the predicted trajectories for BDLO 1 under two manipulation scenarios, using DEFT, a DEFT ablation that leaves out the constraint described in Theorem 4, and Tree-LSTM. The ground-truth initial position of the vertices are colored in blue, the ground-truth final position of the vertices are colored in pink, and the gradient between these two colors is used to denote the ground truth location over time. The predicted vertices are colored as green circles (DEFT), orange circles (DEFT ablation), and light red circles (Tree-LSTM), respectively. A gradient is used for these predictions to depict the evolution of time, starting from dark and going to light. Note that the ground truth is only provided at $t$=0s and prediction is constructed until $t$=8s. The prediction is performed recursively, without requiring additional ground-truth data or perception inputs throughout the entire process.
V-D Compuational Experiment
TABLE III: One Step Inference Time with Dataset 1
Time ($10^{-2}$ s)
Method 1 2 3 4
GCN [33] 0.20 0.19 0.13 0.13
Tree-LSTM [38] 1.81 2.44 1.79 1.89
DEFT 0.76 0.70 0.91 0.93
This section evaluates the computational efficiency of DEFT when compared to baseline methods. It also examines how employing the analytical gradient derived from Theorem 1 to compute (10) and the parallel programming approach detailed in Section IV-F, impacts overall computational time. All experiments in this section were conducted in Python on an Ubuntu 20.04 equipped with an AMD Ryzen PRO 5995WX CPU, 256GB of RAM, and 128 cores.
V-D1 Computational Speed Comparison
We compare computational speed for one step inference. Table III indicates that GCN has the fastest speed. DEFT is the second fastest and is able to generate single step predictions at a 100Hz frequency.
V-D2 Computational Speed with Analytical Gradient
We compare the computational speed of using analytical gradients when compared to using numerical gradients. As summarized in Table IV, computing (10) with the analytical gradient is approximately twice as fast as using numerical gradients.
TABLE IV: Compuational Time for (10) with Dataset 1
Time ($10^{-2}$ s)
Method 1 2 3 4
Analytical Gradient 0.34 0.27 0.41 0.46
Numerical Gradient 0.74 0.65 0.84 0.91
Figure 8: Visualization of planning for BDLO 1 for two manipulation tasks, using DEFT, a DEFT ablation that leaves out the constraint described in Theorem 4, and Tree-LSTM. The BDLO’s goal configuration is highlighted in yellow, while the target hole is shown in red. The DEFT model enables the planning algorithm to successfully complete the task, whereas the ablation approach of DEFT and the Tree-LSTM model both fail to finish tasks.
V-D3 Computational Speed with Parallel Programming
We compare the computational performance of using parallel programming approach described in Section IV-F to a naive (sequential) implementation, in which each branch is simulated one after the other. In this experiment, we use the material properties of BDLO1 and gradually increase the number of children branches from 1 to 8. As shown in Figure 9, the naive implementation’s computation time grows more rapidly than that of the parallel approach. However, a certain increase in computational time is still expected with parallelization. As illustrated in Figure 3, constraints at each junction between the parent branch and its child branches prevent a fully batch-based operation, thereby resulting in an increase in the computational time.
TABLE V: Ablation Study with Dataset 1
Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
Dataset 1 Dataset 2
Method 1 2 3 4 1 2
W/O Constraints.2, 3 and 4 $7.66\times 10^{8}$ $6.10\times 10^{8}$ $6.56\times 10^{6}$ $1.77\times 10^{8}$ $3.81\times 10^{7}$ $1.23\times 10^{8}$
W/O Constraints.3 and 4 $1.36\times 10^{2}$ $1.81\times 10^{2}$ $1.51\times 10^{2}$ $7.64\times 10^{1}$ $8.72\times 10^{1}$ $3.38\times 10^{2}$
W/O Constraint.4 3.95 3.49 2.03 2.15 4.01 4.88
W/O Learning + System ID 2.46 3.53 1.85 1.77 2.03 2.21
W/O System ID 2.24 3.00 1.67 1.50 1.66 1.95
W/O Residual Learning 1.93 3.19 1.70 1.62 1.86 1.89
Full Model 1.87 2.82 1.51 1.41 1.49 1.80
Figure 9: Computational speed comparison: parallel programming versus a naive (sequential) implementation.
V-E Ablation Study
The results presented in Table V demonstrate the contribution of each DEFT component. Notably, excluding Theorems 2, 3, and 4 causes a significant increase in prediction loss. This outcome is primarily driven by the absence of Theorem 2, which leads to simulation instability. Although residual learning is used to mitigate errors, directly modeling stiff behaviors, such as the inextensibility of DLOs, remains challenging, making the learning process highly sensitive to inputs and hampering effective gradient propagation. A key challenge in using neural networks to learn position-based constraints is that constraints such as inextensibility confine the system to a low-dimensional manifold within a high-dimensional space. Without these constraints, it becomes difficult for the network alone to keep the system on that manifold. Over multiple open-loop predictions, even minor errors can accumulate and push the system state off the manifold, causing reduce of residual learning’s effect. Similarly, removing Theorem 3 allows the children branch to detach from parent branch, and therefore fall freely under gravity, leading to a substantial increase in loss. Moreover, excluding Theorem 4 reduces accuracy, demonstrating the importance of introducing rigidity at junctions to allow dynamic propagation across branches.
Despite these challenges, auto-tuning the material properties provides the most substantial improvement over the baseline, illustrating DEFT’s ability to leverage real-world data for parameter identification. Residual learning also yields a non-trivial performance gain. By contrast, the model that do not learn from real-world data underperforms those that do, illustrating the importance of differentiability in this framework.
V-F Planning Results
This section demonstrates how these BDLO models can be used for planning tasks, including shape matching and thread insertion. We compare three approaches: DEFT, which is our proposed method; DEFT without the constraint described in Theorem 4, an ablation approach that removes rigidity at junctions and thus limits dynamic propagation across branches; Tree-LSTM, the baseline model with performance closest to DEFT. In these experiments, we employ ARMOUR [39], a receding-horizon trajectory planner and tracking controller, to manipulate BDLOs. ARMOUR relies on each model’s predicted BDLO states to guide the BDLO to a desired final configuration. Further details on integrating ARMOUR with BDLO simulators are provided in Appendix -E. The results are summarized in Table VI, and Figure 8 shows one of the real-world results. A video of these experiments can be found in the supplementary material.
TABLE VI: Real-World Trajectory Planning Results
Successful Rate
Shape Matching Thread Insertion
Tree-LSTM [38] 2/40 0/35
DEFT w/o Theorem 4 27/40 16/35
DEFT 35/40 29/35
V-F1 Shape Matching
In this experiment a goal configuration for the wire is specified and a random initial condition for the wire is chosen. A trial is considered successful if the Euclidean distance between each ground-truth vertex and its corresponding target vertex after planning is less than 0.05 m. If the task is not completed within 30 seconds, the trial is marked as a failure. As summarized in Table  VI (left column), DEFT achieves a higher success rate than the other two models. Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
V-F2 Thread Insertion
In this experiment, the planning algorithm must insert the end of the BDLO into a hole of radius $2.5$cm which is placed in arbitrary location. The process consists of two stages—shape matching and thread insertion—which together require two goal configurations for the wire. Unlike the previous experiment, in which the robot grasps the branch end, the robot’s end effector clamps the third vertex from the branch’s insertion point. The trial is deemed successful if the BDLO branch successfully passes through the torus while also matching the user-specified wire shape successfully. As shown in Table  VI (right column), DEFT again achieves a higher success rate than the other two models. Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
VI Limitations
While DEFT shows promising results in modeling and control of BDLOs, several limitations remain that could be addressed in future work. First, DEFT currently lacks a contact model for accurately predicting interactions between deformable components and their surroundings. Incorporating a contact modeling framework, such as that proposed in [40], would enable more realistic simulations and more robust planning in environments where collisions and frictional forces play a significant role. Second, for the insertion tasks considered in this work, we assume that the location and geometry of the target hole are known in advance. In unstructured or partially observed environments, this assumption may not hold. Incorporating 3D scene understanding methods, such as Gaussian Splatting [41], could help autonomously detect or refine the location of target features and thereby increasing the system’s applicability to real-world scenarios. Lastly, this work does not account for the orientation of the gripper during BDLO manipulation, which can significantly affect deformation behavior. As discussed in [16], gripper orientation can influence the kinematics of deformable objects. Future extensions of DEFT should therefore incorporate a more comprehensive treatment of gripper orientation to improve accuracy and expand the range of manipulable tasks.
VII Conclusion
This paper introduces DEFT, a method that embeds residual learning within a novel differentiable branched Deformable Linear Object (BDLO) simulator. DEFT enables accurate modeling and real-time prediction of a BDLO’s dynamic behavior over long time horizons under dynamic manipulation. To demonstrate DEFT’s efficacy, we conduct a comprehensive set of experiments evaluating its accuracy, computational speed, impacts of its constribution and utility for manipulation tasks. Compared to the state of the art, DEFT achieves higher accuracy while remaining sample-efficient and without compromising computational speed. This paper also illustrates how to integrate DEFT with a planning and control framework for 3D shape matching and thread insertion of a BDLO in real-world scenarios. DEFT successfully completes more complex, real-world manipulation tasks than other tested baselines.
Acknowledgement
The authors would like to gratefully thank the support by Ford Motor Company.

### v5

@@
 Autonomous wire harness assembly requires robots to manipulate complex branched cables with high precision and reliability.
-A key challenge in automating this process is predicting how these flexible and branched structures behave under manipulation.
-Refer to caption Figure 1: This paper introduces DEFT, a novel framework that combines a differentiable physics-based model with a learning-based approach to accurately model and predict the dynamic behavior of Branched Deformable Linear Objects (BDLOs) in real time.
-As this paper illustrates, this model can be used in concert with a motion planning algorithm to autonomously manipulate BDLOs.
-The figures above illustrate how DEFT can be used to autonomously complete a wire insertion task.
-Top: The system first plans a shape-matching motion, transitioning the BDLO from its initial configuration to the target shape (contoured with yellow), which serves as an intermediate waypoint.
-Bottom: Starting from the intermediate configuration, the system performs thread insertion, guiding the BDLO into the target hole while also matching the target shape.
-Notably, DEFT predicts the state of the wire recursively without relying on ground truth or perception data at any point in the process.
-Without accurate predictions, it is difficult for robots to reliably plan or execute assembly operations.
+A key challenge in automating this process is accurately predicting how flexible branched structures behave during manipulation.
+Without accurate predictions, robots cannot reliably execute assembly operations.
 While existing research has made progress in modeling single-threaded Deformable Linear Objects (DLOs), extending these approaches to Branched Deformable Linear Objects (BDLOs) presents fundamental challenges.
 The junction points in BDLOs create complex force interactions and strain propagation patterns that cannot be adequately captured by simply connecting multiple single-DLO models.
-To address these challenges, this paper presents Differentiable discrete branched Elastic rods for modeling Furcated DLOs in real-Time (DEFT), a novel framework that combines a differentiable physics-based model with a learning framework to: 1) accurately model BDLO dynamics, including dynamic propagation at junction points and grasping in the middle of a BDLO, 2) achieve efficient computation for real-time inference, and 3) enable planning to demonstrate dexterous BDLO manipulation.
-A comprehensive series of real-world experiments demonstrates DEFT’s efficacy in terms of accuracy, computational speed, and generalizability compared to state-of-the-art alternatives.
-I Introduction
+To address these challenges, this paper introduces Differentiable discrete branched Elastic rods for modeling Furcated DLOs in real-Time (DEFT), a novel framework that combines a differentiable physics-based model with a learning framework to: 1) accurately model BDLO dynamics, including dynamic propagation at junction points and grasping in the middle of a BDLO, 2) achieve efficient computation for real-time inference, and 3) enable planning to demonstrate dexterous BDLO manipulation.
+A comprehensive series of real-world experiments demonstrate that DEFT outperforms state-of-the-art methods in accuracy, computational speed, and planning success rate.
+Keywords: Branched Deformable Linear Objects Modeling, Physics-Informed Learning, Differentiable Simulation
+Figure 1: As this paper illustrates, DEFT can be used in concert with a motion planning algorithm to autonomously manipulate BDLOs.
+The figures above illustrate how DEFT can be used to autonomously complete a wire insertion task.
+Left: The system first plans a shape-matching motion, transitioning the BDLO from its initial configuration to the target shape (contoured with yellow), which serves as an intermediate waypoint.
+Right: Starting from the intermediate configuration, the system performs thread insertion, guiding the BDLO into the target hole while also matching the target shape.
+Notably, DEFT predicts the state of the wire recursively without relying on ground truth or perception data at any point in the process.
+1 Introduction
 The automation of wire harness assembly represents a critical challenge in manufacturing, particularly in the automotive and aerospace industries where complex, branched cable systems are essential components.
 These assemblies require robots to manipulate Branched Deformable Linear Objects (BDLOs) with precision over extended periods ($\geq$5s) [1, 2, 3], a task that remains largely unautomated due to the objects’ complex nonlinear dynamics.
+Further complicating matters, practical assembly environments often involve occlusions, limiting perception systems’ ability to estimate full BDLO configurations reliably [4].
+Thus, robust assembly requires accurate BDLO models for reliable long-term open-loop predictions.
 While accurate prediction of BDLO behavior is possible through computationally intensive models, real-time manipulation demands rapid computation.
-Further complicating matters, practical assembly environments often involve occlusions that prevent perception systems from reliably estimating the full BDLO configuration.
-Existing approaches to modeling BDLOs  [4, 5, 6, 7] have focused primarily on computer graphics applications, prioritizing visual plausibility over physical accuracy.
-These models typically require extensive manual parameter tuning and make impractical assumptions about perfect state observation—limitations that render them unsuitable for robotic manipulation tasks.
-Recent advances in modeling single Deformable Linear Objects (DLOs) [8, 9, 10, 11, 12] have shown promise, but the extension to branched structures introduces fundamental new challenges.
-When a robot manipulates one branch of a wire harness, the forces and deformations propagate through junction points to affect all connected branches.
-This coupling between branches, combined with the need for flexible grasp locations and real-time computation, makes BDLO manipulation significantly more complex than handling single DLOs.
-To address these challenges, this paper introduces Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time (DEFT).
-This framework introduces a novel junction coupling approach that accurately models dynamic propagation between parent and child branches while maintaining real-time performance.
-As this paper illustrates through real-world evaluations, DEFT enables three essential wire harness assembly tasks: 3D shape matching, multi-thread grasp, and precise thread grasping and insertion.
-The contributions of this paper are four-fold:
-  1. 1.
-    A novel differentiable BDLO model that accurately captures how dynamics propagate through wire harness junction points thereby enabling model parameters to be efficiently identified via experimental observation;
-  2. 2.
-    A computational representation that enables efficient inference and model parameter learning through parallel programming in concert with analytical gradients;
-  3. 3.
-    A specialized Graph Neural Network (GNN) architecture for BDLO residual learning, which can compensate for numerical integration errors that arise over long time horizon prediction;
-  4. 4.
-    A comprehensive experimental validation demonstrating DEFT’s superior performance in accuracy, speed, and sampling efficiency compared to existing approaches, including integration into a motion planning framework for real-world manipulation tasks.
-The framework is implemented in PyTorch for seamless integration with deep learning workflows, and represents, to the best of our knowledge, the first quantitative investigation of BDLO modeling and manipulation in real-world settings.
-The results demonstrate that DEFT enables reliable autonomous manipulation of complex wire harnesses, addressing a significant challenge in manufacturing automation.
-The code and dataset associated with this paper can be found at https://github.com/roahmlab/DEFT.
-II Related Work
-Effective BDLO modeling requires three critical capabilities: real-time prediction, accurate 3D simulation, and efficient model parameter learning.
-Real-time prediction capability is essential for robotic manipulation tasks, where quick decision-making can mean the difference between successful and failed object handling.
-Three-dimensional accuracy ensures the model can capture the complex spatial deformations that BDLOs undergo during manipulation, including twisting, bending, and branch interactions.
-Efficient parameter learning enables models to automatically adapt to different materials and configurations through data-driven optimization, reducing both computational overhead and the need for extensive training data.
-This section reviews relevant approaches in modeling BDLOs and single DLOs while considering each of these criteria.
-II-A DLOs Modeling
-Approaches to modeling DLOs have evolved from purely physics-based methods toward more sophisticated learning-based frameworks that balance between the three aforemention key capabilities.
-Physics-based approaches implement different mathematical frameworks to model physical behavior.
-Mass-spring systems represent objects as networks of point masses connected by springs, computing forces and accelerations to simulate movement [13, 14, 15].
-Position-Based Dynamics (PBD) directly manipulates vertex positions to satisfy physical constraints, avoiding expensive force calculations [10, 16, 9].
-Finite Element Methods (FEM) divide objects into discrete elements and solve partial differential equations to compute deformations [17, 18, 19].
-Discrete Elastic Rods (DER) models represent objects as centerline curves with material frames, enabling explicit modeling of bending and twisting [16, 20, 21, 22].
-Unfortunately each of these methods has limitations: mass-spring systems describe the rigidity of wires by introducing large spring stiffness which can make prediction numerically unstable; PBD models struggle with parameter learning because they assume that the system is quasi-static; FEM methods are computationally intensive; and DER models lack differentiable expressions for their model parameters, making efficient parameter optimization challenging.
-Learning-based modeling approaches use data-driven techniques to predict deformation behavior.
-Methods such as Graph Neural Networks (GNN) process vertex-level spatial information [12], while Bi-LSTM models capture temporal dependencies in deformation sequences [11].
-These approaches excel at real-time prediction through efficient computation and support straightforward parameter learning through standard optimization techniques.
-However, they struggle with 3D accuracy, particularly in capturing the complex dynamic deformations characteristic of real-world manipulation.
-Recent work has explored hybrid approaches that combine learning with physics-based modeling.
-This is exemplified by DEFORM [8] which constructs a novel differentiable DER model and combines it with a neural network-based residual learning approach to account for numerical integration errors during prediction.
-This hybrid approach achieves both real-time performance and 3D accuracy while enabling efficient parameter learning through end-to-end differentiability.
-Despite DEFORM’s advances in DLO modeling, it faces three significant limitations when extended to BDLO applications.
-First, DEFORM’s modeling framework does not address the dynamics of branched structures, particularly the propagation of stretching, twisting, and bending forces at junction points.
-Second, DEFORM’s manipulation framework assumes fixed clamping points at the object ends, following conventional DLO literature.
-This assumption restricts its applicability to more sophisticated manipulation tasks, such as thread insertion, which require grasping at multiple points along the object’s length.
-Third, DEFORM’s computational approach presents efficiency challenges for BDLO simulation.
-Its sequential processing of individual branches would result in linear computational time increases as branches are added.
-Furthermore, its bi-level optimization framework, while enabling learning from real-world data, relies on numerical gradients that can lead to slow convergence and optimization instability.
-II-B BDLOs Modeling
-Existing BDLO modeling approaches adapt various physical simulation frameworks to handle branched structures.
-FEM-based methods extend traditional finite element analysis by treating branches as connected elastic elements [23].
-While these methods excel in 3D accuracy and enable efficient parameter learning through gradient-based optimization, their computational requirements preclude real-time prediction.
-DER-based approaches introduce rigid-body constraints at junctions to ensure proper alignment of attached segments  [4], achieving real-time performance and reasonable 3D accuracy but struggling with parameter learning due to their fixed parameter structure.
-Quasi-static methods like ASMC enforce geometric constraints while assuming static equilibrium conditions [5], offering real-time performance, but sacrificing 3D accuracy by overlooking dynamic effects that can make parameter learning challenging.
-To address these challenges, DEFT introduces a novel differentiable framework specifically designed for BDLOs.
-Through a specialized GNN architecture for branched structures, DEFT achieves all three critical capabilities: real-time performance, 3D accuracy, and efficient parameter learning.
-This enables DEFT to adapt to various materials and configurations while maintaining the speed and accuracy necessary for practical robotic manipulation tasks.
-III Preliminaries
-This section introduces the key components involved in modeling DLOs. It then summarizes the primary contributions of DEFORM and describes the challenges of adapting it to BDLO modeling.
-III-A Deformable Linear Object (DLO) Model
-To model a DLO, we represent it as an indexed set of $n$ vertices at each time instance.
+Deep neural networks (DNNs) for tree-structured data [5, 6, 7, 8] can achieve real-time prediction performance for BDLOs but often require large datasets and lack robustness to variations in BDLOs.
+A promising alternative combines differentiable physics models as high-fidelity priors with DNNs to improve sample efficiency and generalizability [9, 10, 11, 12, 13, 14, 15].
+However, this approach has not yet been applied to BDLO dynamics due to two significant challenges: First, the field lacks an accurate, efficient, differentiable physics model for BDLOs.
+Current models compromise speed for accuracy and do not preserve gradients needed for parameter optimization.
+Second, to the best of our knowledge, there is no existing research demonstrating that DNNs can effectively learn residual dynamics specifically for BDLOs.
+To address these challenges, this paper introduces Differentiable Branched Discrete Elastic Rods for Modeling Furcated DLOs in Real-Time (DEFT), a novel differentiable physics-based framework that accurately predicts dynamic BDLO behavior over long time horizons with real-time inference capabilities.
+The contributions of this paper are four-fold: First, a novel differentiable BDLO model that accurately captures dynamics propagation at junctions, enabling efficient parameter identification.
+Second, a computational representation that enables efficient inference and model parameter learning through parallel programming combined with analytical gradients.
+Third, a specialized Graph Neural Network (GNN) architecture for residual learning to correct numerical integration errors over long horizons.
+Fourth, a comprehensive experimental validation demonstrating that DEFT outperforms prior methods in accuracy, speed, and sample efficiency while enabling practical applications such as 3‑D shape matching, multi‑branch grasping, and thread insertion (Figure 1).
+The framework is implemented in PyTorch for seamless integration with deep learning workflows, and represents, to the best of our knowledge, the first quantitative study of BDLO modeling and manipulation in real-world settings.
+A visualization of DEFT’s contributions can be found in Appendix B.1.
+2 Related Work
+Physics-Based Modeling.
+Several methods adapt tree or cloth modeling to branched structures.
+FEM-based approaches [16, 17, 18, 19, 20, 21, 22] offer high 3D accuracy but require extensive parameter tuning and are computationally intensive.
+Discrete Elastic Rods (DER) [23, 24] model branches individually using elastic rod theory with rigid-body constraints at junction points.
+These methods are reasonably accurate, but rely on fixed parameters that limit learning and require precise orientation tracking, which is impractical for robotic tasks.
+Quasi-static methods like ASMC [25] prioritize speed by enforcing geometric constraints, which can make parameter identification difficult and compromise dynamic fidelity.
+Learning-Based Modeling.
+Unlike physics-based approaches, neural networks learn from real-world data.
+While Tree-LSTM [5, 26] and Graph Neural Networks (GNNs) [6, 8, 7] can process vertex-level spatial information for BDLOs, our experiments show they struggle to accurately capture complex 3D dynamic deformations during actual manipulation tasks.
+Physics-Learning Hybrid Modeling.
+Recent research combines physics-based modeling with learning [10, 11, 12, 13, 14, 15].
+For instance, DEFORM [14] focuses on modeling single-threaded DLOs using a differentiable DER (DDER) model and residual learning to correct numerical integration errors.
+However, DEFORM lacks fundamental capabilities essential for accurate and efficient BDLO modeling.
+Our experiments show that extending DEFORM to BDLOs reduces computational efficiency and fails to capture branching point dynamics, leading to inaccurate predictions.
+DEFT addresses these limitations by adapting DDER for individual branches and introduces several novel contributions (as described in the Introduction) for accurate, efficient BDLO simulation.
+Comprehensive ablation studies and comparisons validate DEFT’s advantages over DEFORM extensions and state-of-the-art methods.
+3 Preliminaries
+Deformable Linear Object (DLO) Model.
+We represent a DLO as an indexed set of $n$ vertices at each time instance.
 The ground truth location of vertex $i$ at time $t$ is denoted by $\mathbf{x}_{t}^{i}\in\mathbb{R}^{3}$, and the set of all $n$ vertices is denoted by $\mathbf{X}_{t}$.
@@
 We distinguish between ground truth elements and predicted elements by using the circumflex symbol (e.g., $\mathbf{X}_{t}$ is the ground truth set of vertices at time $t$, and $\hat{\mathbf{X}}_{t}$ is the predicted set of vertices at time $t$).
-III-B Modeling DLOs with DEFORM
-III-B1 Discrete Elatic Rods (DER)
-DEFORM builds upon DER theory to model three distinct deformation modes of DLOs: bending, twisting, and stretching.
-DER theory models DLOs using two key families of coordinate frames.
+Differentiable Discrete Elastic Rods (DDER).
+DDER builds upon DER theory to model three distinct deformation modes of DLOs: bending, twisting, and stretching while being able to auto-tune model parameters from real-world dataset.
+We briefly summarize DER in this section.
+A longer introduction can be found in Appendix A).
+To predict $\hat{\mathbf{X}}_{t+1}$, DER assigns a scalar angle $\theta^{\text{i}}_{t}\in\mathbb{R}$ to each segment and uses these angles to construct $n-1$ material frames at time $t$, denoted by $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$.
+Let $\bm{\alpha}$ represent the vector of material properties for each vertex of the DLO (i.e., mass, bending modulus, and twisting modulus) [23, §4.2].
+Using these definitions, one can compute the potential energy of the DLO arising from bending and twisting, which we denote by $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$.
+At each step, DER assumes quasi‑static equilibrium, calculates the optimal angles $\bm{\theta}_{t}^{*}$ by minizing $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to the decision variables $\bm{\theta}_{t}$.
+Once $\bm{\theta}^{*}_{t}(\mathbf{X}_{t},\bm{\alpha})$ is derived, the restorative force during deformation is given by the negative gradient of the potential energy with respect to the vertices.
+Consequently, the equation of motion for the DLO is:
+$\mathbf{M}\ddot{\textbf{X}}_{t}=-\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}}$ (1)
+One can numerically integrate this formula to predict the velocity and position $\hat{\textbf{X}}_{t+1}$ by applying the Semi-Implicit Euler method.
+Consequently, minimizing the prediction loss requires solving a bi-level optimization problem at each time step $t$:
+$\underset{\bm{\alpha}}{\min}\;\bigl{\lVert}\mathbf{X}_{t+1}-\hat{\mathbf{X}}_{t+1}(\bm{\theta}_{t}^{*})\bigr{\rVert}_{2}$ (2)
+$\hskip 56.9055pt\bm{\theta}_{t}^{*}=\underset{\bm{\theta}_{t}}{\arg\min}\;P\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}$ (3)
+Note $\hat{\textbf{X}}_{t+1}$ depends on the previously computed $\bm{\theta}^{*}_{t}$ through (3).
+To compute the gradient of (3) with respect to $\bm{\alpha}$, DDER employs solvers from Theseus [27], which use PyTorch’s automatic differentiation [28].
+As we illustrate via experiments in §5, this can be a time-consuming process.
+4 Methodology
+4.1 Branched Deformable Linear Object (BDLO) Model
+Figure 2: Left: An example configuration of BDLO manipulation, relevant notation and structure at a junction.
+Right: The adjacency matrix $\mathbf{A}$ for the junction in left captures both self-loops (diagonal entries) and inter-node connections (off-diagonal entries).
+By embedding each node’s local dynamics alongside its coupling to neighboring nodes, $\mathbf{A}$ enables an graph representation of the BDLO’s behavior.
+We model a BDLO by decomposing its structure into a hierarchy of DLOs.
+The parent DLO is defined as the minimum-length path through the BDLO that connects the two end-effectors (i.e., the points where the structure is manipulated or anchored).
+All other branches diverging from this primary path are modeled as children DLOs that maintain their physical connection to the parent DLO at their attachment points, which we refer to as junction points.
+The parent and children DLOs are each modeled as DLOs using DDER and inherit all previously introduced notation.
+Denote the parent and child DLOs using superscripts $p$ and $c$, respectively.
+At a junction located at vertex $i$, associated edges are $\mathbf{e}^{p,i-1}_{t}$, $\mathbf{e}^{p,i}_{t}$, and $\mathbf{e}^{c,1}_{t}$ as shown in Figure 2 (left).
+To capture bending and twisting effects at the junction, we introduce the moment of inertia $\mathbf{I}^{\text{i}}\in\mathbb{R}^{3\times 3}$ for the $i$-th edge, representing mass distribution resistance to rotation.
+Edge orientation is described by an angle-axis vector $\mathbf{\Omega}^{i}_{t}\in\mathbb{R}^{3}$, defined in Appendix B.2.
+The orientation and position of the two end effectors at time $t$ are denoted by $\mathbf{u}_{t}\in\mathbb{R}^{12}$.
+$\mathbf{u}_{t}$ is referred to as the input.
+4.2 DEFT Overview
+Algorithm 1 $(\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}})=\textbf{DEFT}(\hat{\mathbf{X}}_{\text{t}},\hat{\mathbf{V}}_{\text{t}},\textbf{u}_{t}$)
+1:Associate grasped vertices with $\textbf{u}_{t}$
+2:Calculate Material Frames $\Gamma(\textbf{X}_{t},\bm{\theta}_{t})$
+3:$\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$triangleright §4.5
+4:$\hat{\textbf{V}}_{t+1}=\hat{\textbf{V}}_{t}-\Delta_{t}\mathbf{M}^{-1}\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}}$triangleright (1)
+5:$\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t}+\Delta_{t}\hat{\textbf{V}}_{t+1}+\text{GNN}$triangleright §4.3
+6:Constraints enforcement on $\hat{\mathbf{X}}_{t+1}$triangleright §4.4
+7:$\hat{\textbf{V}}_{t+1}=(\hat{\mathbf{X}}_{t+1}-\hat{\mathbf{X}}_{t})/\Delta_{t}$triangleright Velocity Update
+return $\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}}$
+At each time $t$, DEFT simulates the behavior of a BDLO when given predictions of the vertices and their velocities under a specified input as described in Algorithm 1.
+It begins by first associating the grasped vertices with the robot end effector inputs $\mathbf{u}_{t}$.
+DEFT then computes their respective material frames, which are subsequently used to calculate each branch’s potential energy (Line 2).
+DEFT then minimizes this potential energy using analytical gradients to determine the optimal parameters $\bm{\theta}_{t}^{*}$ and updates the potential energy accordingly (Line 3).
+These forces are integrated forward in time using a Semi-Implicit Euler method combined with residual learning to update the vertex velocities and positions (Lines 4 and 5).
+To further accelerate above computation, DEFT processes the parent and child DLOs in parallel.
+Numerical integration errors can cause unintended changes in branch lengths.
+Additionally, DEFT initially treats each branch’s internal dynamics independently, ignoring junction interactions, which can impede the propagation of dynamics, such as bending and twisting, across junctions or even cause branch detachment.
+To overcome this, DEFT enforces constraints ensuring branch inextensibility, junction connectivity, and proper propagation of deformation effects (Line 6).
+Applying Algorithm 1 iteratively enables accurate BDLO state predictions over extended time horizons.
+An algorithm overview is illustrated in Appendix B.1.
+4.3 Integration Method with Residual Learning
+This subsection constructs the GNN for residual learning in integration method.
+We employ a GNN to represent BDLOs as graphs where vertices become nodes and edges capture both intra-branch connections and junctions.
+This graph-based formulation translates into an adjacency matrix $\mathbf{A}\in\mathbb{R}^{(n^{p}+n^{c})\times(n^{p}+n^{c})}$, which links individual branch nodes and explicitly encodes node information across branch junctions.
+This adjacency matrix is defined as follows: the entry $\mathbf{A}_{i_{1},i_{2}}$ equals $1$ if there’s an edge between vertices $i_{1}$ and $i_{2}$, and 00 otherwise.
+Figure 2 depicts an example $\mathbf{A}$.
+This adjacency matrix is utilized within a Graph Convolutional Network (GCN) [8] to aggregate learned features.
+A single GCN layer is defined as
+$\text{GCN}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})=\mathbf{\tilde{A}}\,\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\,\mathbf{W},\hskip 5.69054pt\mathbf{\tilde{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (4)
+where $\mathbf{\tilde{A}}$ is the normalized adjacency matrix.
+Note that $\mathbf{\tilde{A}}$ is the normalization of $\mathbf{A}$ using the degree matrix $\mathbf{D}=\mathrm{diag}(d_{1},d_{2},\ldots,d_{n})$ where $d_{i1}=\sum_{i2}A_{i1,i2}$.
+This normalization ensures nodes with higher degrees do not disproportionately influence the model, maintaining consistent feature magnitudes across representations [29].
+The details node feature matrix, $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\in\mathbb{R}^{(n^{p}+n^{c})\times h}$ and the trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ can be found in Appendix B.3.
+4.4 Enforcing Constraints
+After numerical integration, DEFT applies constraints to maintain branch inextensibility, junction attachment, and enable deformation propagation by updating vertex positions using correction terms.
+Direct projection onto constraint sets may cause nonlinear momentum changes, and can lead to simulation instability and inaccurate system identification.
+Alternatively, inspired by the PBD method [30], DEFT computes correction terms via optimization to preserve momentum, resulting in stable and realistic simulations.
+Below, we introduce a theorem that can be applied to enforce inextensibility, attachment at the junctions, and deformation propagation.
+Note that we omit the time subscript $t$ in the statement of the theorem for simplicity.
+Theorem 1.
+Consider two consecutive bodies a and b, represented by adjacent vertices or edges, that are subject to a holonomic constraint $C(\hat{\mathbf{z}})=0$, where $\hat{\mathbf{z}}=\bigl{[}\hat{\mathbf{x}}^{a},\,\hat{\mathbf{x}}^{b},\,\hat{\bm{\Omega}}^{a},\,\hat{\bm{\Omega}}^{b}]$.
+Suppose $\hat{z}$ does not satisfy the holonomic constraint and one calculates a correction $\Delta\hat{\mathbf{z}}$ to $\hat{z}$ to satisfy the holonomic constraint by solving the following momentum constrained optimization problem.
+$\min_{\Delta\hat{\mathbf{z}}}\;\tfrac{1}{2}\,C\bigl{(}\hat{\mathbf{z}}+\Delta\hat{\mathbf{z}}\bigr{)}^{2},$ (5)
+[…]
+[…]
+If $C$ is affine, differentiable, and symmetric in $\Delta\hat{\mathbf{z}}$, then when $\Delta_{t}$ is sufficiently small there exists an analytical formula for the unique solution to the optimization problem.
+The proof of this theorem for each of the relevant constraints and the formula for the optimal solution can be found in Appendix B.4.
+We next demonstrate how this theorem can be applied to enforce inextensibility, junction attachment, and deformation propagation.
+To enforce inextensibility, let $\overline{\mathbf{e}}_{i}$ denote the undeformed edge between vertices $\hat{\mathbf{x}}^{i}$ and $\hat{\mathbf{x}}^{i+1}$.
+Set $\hat{\mathbf{z}}_{I}=\bigl{[}\hat{\mathbf{x}}^{i},\,\hat{\mathbf{x}}^{i+1},\,\hat{\bm{\Omega}}^{i},\,\hat{\bm{\Omega}}^{i+1}],C_{I}(\hat{\mathbf{z}}+\Delta\hat{\mathbf{z}})=||(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})-(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})||_{2}-||\overline{\textbf{e}}_{i}||_{2}$ so that the updated edge length is equal to its undeformed length.
+To enforce attachment at the junction, define $\hat{\mathbf{z}}$ and $g(\hat{\mathbf{z}}+\Delta\hat{\mathbf{z}})$ between parent‐branch vertex $\hat{\mathbf{x}}^{p,i}$ and child‐branch vertex $\hat{\mathbf{x}}^{c,1}$ as $\hat{\mathbf{z}}_{A}=\bigl{[}\hat{\mathbf{x}}^{p,i},\,\hat{\mathbf{x}}^{c,1},\,\hat{\bm{\Omega}}^{p,i},\,\hat{\bm{\Omega}}^{c,1}],C_{A}(\hat{\mathbf{z}}+\Delta\hat{\mathbf{z}})=||(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})-(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})||_{2}$ to enforce zero separation at the junction.
+To enable deformation propagation, a key observation is that many practical wire-harnesses feature junctions that are nearly rigid.
+This near-rigidity arises because manufacturers often use rigid plastics or other stiff materials at these connection points [33].
+To capture this rigidity, we set $\hat{\mathbf{z}}_{O}=\bigl{[}\hat{\mathbf{x}}^{p,i},\,\hat{\mathbf{x}}^{c,1},\,\hat{\bm{\Omega}}^{p,i},\,\hat{\bm{\Omega}}^{c,1}]$ and $C_{O}(\hat{\mathbf{z}},\Delta\hat{\mathbf{z}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{x}},\Delta\hat{\mathbf{x}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{x}},\Delta\hat{\mathbf{x}})\big{)}-\big{(}\hat{\bm{\Omega}}^{c,1}(\hat{\mathbf{x}},\Delta\hat{\mathbf{x}})+\hat{\bm{\Omega}}_{\Delta}^{c,1}(\hat{\mathbf{x}},\Delta\hat{\mathbf{x}}\big{)}\|_{2}-\epsilon,$ where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix B.2 and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid.
+4.5 Algorithm Efficiency Improvement
+Modeling speed is critical for efficient planning.
+However, as demonstrated in the experiments, directly implementing Algorithm 1 results in slow simulation and planning.
+This section presents two methods to significantly enhance computational performance.
+Analytical Gradient Derivation: The goal of this subsection is to present the analytical gradient of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$.
+To improve the paper’s readability, we summarize the result in the following theorem whose proof can be found in Appendix B.5.
+Theorem 2.
+Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ be the Material Frame at time $t$, and let $\bm{\alpha}$ contain the bending stiffness $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and twisting stiffness $\beta\in\mathbb{R}$.
+Suppose the total potential energy is the sum of bending and twisting potential energies.
+The partial derivative of the bending and twisting potential energies with respect to $\theta^{i}$ are given by
+$\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},$ (8)
+$\frac{\partial P_{\mathrm{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\beta^{i}\,\bigl{(}\theta^{i}-\theta^{i-1}\bigr{)}\;-\;\beta^{i+1}\,\bigl{(}\theta^{i+1}-\theta^{i}\bigr{)},$ (9)
+respectively, where $\bm{\omega}^{(i,j)}$ is the material curvature.
+Parallel Programming.
+Unlike quasi-static approaches [24, 25] that constantly require handling interactions between parent and child branches at each time step, DEFT constructs a computational graph that computes dynamics at junctions through constraints as post-processing.
+This strategy enables efficient parallel execution of Algorithm 1 (Lines 1-5 and 7) via PyTorch’s batch-wise operations, seamlessly integrating the physics-based framework with residual learning methods.
+To address dimensional mismatches between parent ($n^{p}$) and child ($n^{c}$) branches, we pad child branch matrices and vertices with zeros, ensuring uniform batch dimensions.
+Additionally, gradients from padded elements are masked during parameter updates to avoid influencing learned parameters.
+5 Experiments and Results
+Figure 3: Visualization of predicted trajectories for BDLO 1 under scenarios where one robot grasps an end and the other grasps the midpoint, comparing DEFT, a DEFT ablation without enforcing junction near-rigidity (§4.4), and Tree-LSTM.
+Ground-truth vertex positions transition from blue (initial) to pink (final) via gradients indicating intermediate positions.
+Predictions similarly use gradients from dark to light.
+Ground truth is provided only at $t=0$ s, with recursive predictions performed until $t=8$ s without additional perception inputs.
+Dataset Collection and Training.
+We collect dynamic trajectories for selected BDLOs using a motion capture system operating at 100 Hz, capturing both slow and fast motions.
+Data collection was divided into two categories: Dataset 1: Both robots grasped BDLO ends (for all BDLOs).
+Dataset 2: One robot grasped an end, and the other grasped the midpoint (for BDLOs 1 and 3).
+Experimental setup and data collection details are provided in Appendix C.1.
+Each method’s performance is evaluated when predicting the state of the wire over a 500 step (i.e.
+5s) prediction horizon using just the state of the wire at $t=0$ s and the orientation and position of the end effectors at each time step.
+Table 1: Left: Modeling Results Over a $5$s Horizon. Right: Average Computational Time per 0.01 s Inference.
+Modeling Accuracy (RMSE, $10^{-2}$ m) Computational
+Dataset 1 Dataset 2 Time ($10^{-2}$s)
+Method 1 2 3 4 1 3 1
+GCN [8] 6.70 6.42 5.25 4.20 7.33 4.24 0.163
+Tree-LSTM [31] 5.32 4.50 2.46 2.71 7.08 4.55 1.98
+FEM [32]: $\Delta_{t}$: $10^{-6}$s, node: 60 3.11 3.70 2.29 2.84 2.17 2.10 493
+FEM [32]: $\Delta_{t}$: $10^{-6}$s, node: 160 2.07 3.35 1.82 1.54 1.40 1.94 1211
+FEM [32]: $\Delta_{t}$: $10^{-2}$s, node: 60 14.6 18.3 12.1 17.6 14.9 12.3 0.485
+FEM [32]: $\Delta_{t}$: $10^{-2}$s, node: 160 12.9 15.1 10.0 11.8 12.8 9.71 1.09
+DEFORM+[14] 4.59 7.44 5.02 4.46 3.88 4.25 1.665
+DEFT 1.87 2.82 1.51 1.41 1.49 1.80 0.825
+Baseline Comparisons.
+Only a limited number of studies have explored the dynamic behavior of BDLOs, and none provide publicly available code, making direct replication infeasible.
+We select four implementations for comparison.
+Learning-Based Modeling: As BDLOs naturally map onto graph structures, architectures like Tree-LSTM [31] and GCN [8] could model their dynamics.
+FEM-Based Modeling: We selected FEM [32] as a baseline and optimized its parameters via grid search.
+Simulations varied timestep sizes from $1\times 10^{-2}$ s (DEFT’s setting) to $1\times 10^{-6}$ s and used 60 or 160 nodes, compared to DEFT’s $\sim$20.
+DEFORM+: To highlight DEFT’s contributions, we created DEFORM+, which extends DEFORM[14] to BDLOs by replacing DEFT’s junction constraints (§4.4) with FEM-based springs and bending.
+DEFORM+ excludes DEFT’s residual learning architecture(§4.3) and DEFT’s computational optimizations (§4.5).
+This is the most straightforward extension of DEFORM to BDLOs.
+Modeling Results Table 1 (left) summarizes the average RMSE comparing DEFT to baselines across two dataset categories over a 5-second prediction horizon.
+DEFT outperforms all baselines in all but one case, where it is outperformed by the finest FEM.
+Figure 3 visualizes performance comparisons among DEFT, Tree-LSTM, and a DEFT ablation described later.
+Experimental videos are available in the supplementary material.
+Table 1 (right) summarizes average computational time for 0.01s inference.
+Table 1 indicates that GCN has the fastest speed.
+DEFT is the third fastest and is able to generate single step predictions above 100Hz frequency.
+In the planning task described later, DEFT completes each optimization loop 8.8 times faster than DEFORM++ and 1952 times faster than the finest FEM baseline.
+More experiments of computational speed are in Appendix C.3.
+Ablation Study Table 2 highlights the impact of each DEFT contribution.
+Excluding all constraints introduced in §4.4 results in simulation instability.
+This outcome is primarily driven by the absence of inextensibility enforcement, which leads to simulation instability.
+Similarly, removing attachment constraints allows the children branch to detach from parent branch and fall freely, leading to a substantial increase in loss.
+Excluding near-rigidity constraints similarly reduces accuracy, underscoring the necessity of junction rigidity for effective dynamic propagation across branches.
+Note Although residual learning is used to mitigate errors, directly modeling stiff behaviors as described in §4.4, remains challenging, making the learning process highly sensitive to inputs and hampering effective gradient propagation.
+Neural networks alone struggle to maintain the system within the low-dimensional manifold defined by these constraints and minor errors can accumulate during open-loop predictions, reducing residual learning effectiveness.
+Despite these challenges, combination of auto-tuning the material properties and residual learning also provides the substantial improvement, illustrating DEFT’s ability to leverage real-world data for parameter identification and residual learning.
+By contrast, the model that do not learn from real-world data underperforms those that do, illustrating the importance of differentiability in this framework.
+Table 2: Ablation Study with Dataset 1
+Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
+Dataset 1
+Method 1 2 3 4
+W/O Inext, Attach & Near-Rigidity $7.66\times 10^{8}$ $6.10\times 10^{8}$ $6.56\times 10^{6}$ $1.77\times 10^{8}$
+W/O Attach & Near-Rigidity $1.36\times 10^{2}$ $1.81\times 10^{2}$ $1.51\times 10^{2}$ $7.64\times 10^{1}$
+W/O Near-Nigidity 3.95 3.49 2.03 2.15
+W/O Learning + System ID 2.46 3.53 1.85 1.77
+W/O System ID 2.24 3.00 1.67 1.50
+W/O Residual Learning 1.93 3.19 1.70 1.62
+Full Model 1.87 2.82 1.51 1.41
+Figure 4: Visualization of planning for BDLO 1 for thread insertion, using DEFT, DEFT ablation without enforcing junction near-rigidity (§4.4), and Tree-LSTM.
+The BDLO’s goal configuration is highlighted in yellow, while the target hole is shown in red.
+The DEFT model enables the planning algorithm to successfully complete the task, whereas the ablation approach of DEFT and the Tree-LSTM model both fail to finish tasks.
+Shape Matching.
+In these experiments, we use BDLO 1 and employ ARMOUR [33], a receding-horizon trajectory planner and tracking controller, to manipulate the BDLO based on each model’s predictions.
+Integration details of ARMOUR with DEFT are in Appendix D.
+Each trial specifies a goal configuration and a random initial condition.
+A trial succeeds if each vertex reaches within 0.05 m of its target within 30 seconds.
+As shown in Table 3 (left), DEFT outperforms both baselines.
+Thread Insertion.
+In this experiment, the planning algorithm must insert the end of the BDLO into a hole of radius $2.5$ cm which is placed in arbitrary location.
+The process consists of shape matching and thread insertion.
+Note that shape matching prevents deviations from the desired wire shape, avoiding excessive twisting or bending that could compromise safety and subsequent tasks (e.g., wire layout).
+Unlike the previous experiment, in which the robot grasps the branch end,
+Table 3: Real-World Trajectory Planning Results
+Success Rate
+Methods Shape Matching Insertion
+Tree-LSTM [31] 2/40 0/35
+DEFORM+ [14] 10/40 4/40
+DEFT (Ablation) 27/40 16/35
+DEFT 35/40 29/35
+the robot’s end effector clamps the third vertex from the branch’s insertion point.
+The trial is deemed successful if the BDLO branch passes through the torus while matching the user-specified wire shape.
+The results are summarized in Table 3 (right), and Figure 4 shows one of the real-world results.
+6 Conclusion
+This paper introduces DEFT, a method embedding residual learning within a novel differentiable branched Deformable Linear Object (BDLO) simulator.
+DEFT delivers accurate modeling and real-time prediction of BDLO dynamic behavior during manipulation over extended time horizons.
+Compared to state-of-the-art approaches, DEFT achieves superior accuracy while maintaining sample efficiency and computational speed.
+We demonstrate DEFT’s integration with planning and control frameworks for 3D shape matching and thread insertion in real-world scenarios, successfully completing more complex manipulation tasks than other baselines.
+7 Limitations
+While DEFT demonstrates strong performance in modeling and controlling BDLOs, several areas for improvement remain.
+First, DEFT assumes the locations and geometries of the target insertion holes are known.
+This assumption may not hold in unstructured or partially observable environments.
+Incorporating advanced 3D scene understanding techniques, such as Gaussian Splatting [34], could improve system autonomy and adaptability.
+Second, DEFT occasionally fails at insertion tasks due to neglecting the gripper orientation during BDLO manipulation.
+As highlighted in [35], gripper orientation can affect a deformable object’s kinematics.
+Extending DEFT to explicitly model gripper orientation could enhance reliability.
+Lastly, DEFT does not explicitly handle interactions between deformable objects and their environment, such as contact forces and friction.
+Integrating contact modeling frameworks ([36]) would enable more accurate simulations and robust performance in scenarios involving frequent collisions or tight placements.
+Due to space constraints and the focused scope of this work, comprehensively addressing these limitations is deferred to future research.
+8 Acknowledgement
+The authors would like to gratefully thank the support by Ford Motor Company.
+Appendix A Supplementary Material for DER Background
+To model bending, twisting, and stretching, DER theory models DLOs using two key families of coordinate frames.
 First, it introduces Bishop Frames, which provide a twist-free reference state along the centerline of the rod.
@@
 Second, the theory employs Material Frames, which describe the actual physical deformation of the rod.
+Figure 5: An illustration of DER coordinate frames.
 Material Frames can be expressed relative to the Bishop Frames through a single angle whose rate of change represents the physical twist of the DLO.
@@
 An illustration of the frames can be found in Figure 5.
-Figure 2: An illustration of DER coordinate frames.
-Equations of Motion: Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ denote the set of all $n-1$ Material Frames at time $t$.
-Let $\bm{\alpha}$ represent the vector of material properties for each vertex of the DLO (i.e., mass, bending modulus, and twisting modulus) [4, Section 4.2].
-Using these definitions, one can compute the potential energy of the DLO arising from bending and twisting which we denote by $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$.
-DER theory assumes that the DLO reaches an equilibrium state between each simulation time step to obtain the optimal $\bm{\theta}_{t}^{*}$:
+Equations of Motion: DER theory assumes that the DLO reaches an equilibrium state between each simulation time step to obtain the optimal $\bm{\theta}_{t}^{*}$:
 $\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ (14)
@@
 where $\Delta_{\text{t}}>0$ is a user specified time discretization.
-III-B2 Differentiable Discrete Elastic Rods (DDER)
-Tuning the DER model parameters, $\bm{\alpha}$, to match real-world DLOs is non-trivial.
-To address this challenge, DEFORM introduces DDER, which enables the automatic tuning of DLO parameters using real-world data.
-Predicting $\hat{\textbf{X}}_{t+1}$ with DDER through (7) and (8) involves solving an optimization problem as defined in (5).
-Consequently, minimizing the prediction loss requires solving a bi-level optimization problem at each time step $t$:
-$\underset{\bm{\alpha}}{\min}\hskip 14.22636pt$ $\|\textbf{X}_{t+1}-\hat{\textbf{X}}_{t+1}(\bm{\theta}^{*}_{t})\|_{2}\hskip 42.67912pt$ (9)
-$\bm{\theta}^{*}_{t}=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ (10)
-Note $\hat{\textbf{X}}_{t+1}$ depends on the previously computed $\bm{\theta}^{*}_{t}$ through (6), (7) and (8).
-To compute the gradient of (10) with respect to $\bm{\alpha}$, DEFORM employs solvers from Theseus [25], which use PyTorch’s automatic differentiation [26].
-As we illustrate via experiments in Section V, this can be a time-consuming process.
-III-B3 Integration Method with Residual Learning
-To compensate for numerical errors that arise during numerical integration(8), DEFORM incorporates a learning-based framework into the integration method:
-$\hat{\textbf{X}}_{t+1}=\hat{\textbf{X}}_{t}+\Delta_{t}(\hat{\textbf{V}}_{t+1}+\text{GNN}(\hat{\textbf{X}}_{t},\hat{\textbf{V}}_{t},\bm{\alpha}))$ (11)
-This structure builds on residual learning [27], grounding the neural network in physical laws while leveraging data-driven insights.
-III-B4 Inextensiblity Enforcement with Momentum Conservation
-After numerical integration, the DLO length may deviate from its initial value due to numerical errors, making it difficult to accurately model inextensible DLOs.
-A common physics-based approach [13, 14, 15] uses mass-spring systems to minimize this deviation, but this can result in stiff simulations and numerical instability.
-DEFORM addresses this by enforcing an inextensibility constraint while preserving momentum, leading to more accurate and stable simulation.
-III-B5 Challenges of Extending DEFORM to BDLOs
-Though DEFORM can accurately simulate single-thread DLOs, it is unsuitable for BDLOs for several reasons.
-First, the definitions of Bishop and Material Frames do not generalize to BDLOs.
-To see why this is true, recall that at junctions in a BDLO there are several tangent vectors because there are more than two edges.
-As a result, satisfying (1) and (2) results in an overdetermined system which makes it impossible to find a rotation matrix $\mathbf{R}^{i}$ that satisfies all constraints.
-Consequently, the Bishop and Material Frames cannot be defined at these junctions.
-An alternative to extend DEFORM to BDLOs is to simulate each branch individually using DEFORM.
-However, this requires repeatedly solving (5), which reduces computational efficiency.
-This is further complicated when trying to enforce inextensibility using the momentum conservation constraint.
-To address this, DEFT proposes a computational representation that supports parallel processing, enabling more efficient inference.
-Second, relying on automatic differentiation to perform parameter tuning becomes computationally prohibitive when dealing with BDLOs.
-To address this limitation, DEFT utilizes the analytical gradients of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$.
-Third, the GNN used by DEFORM during residual learning is tailored to single-strand DLOs.
-To address this limitation, DEFT introduces a specialized GNN Architecture for residual learning for BDLOs that can compensate for numerical integration errors over long horizon predictions.
-Figure 3: Algorithm Overview of DEFT.
-In the initialization stage (Section IV-A), DEFT begins by separating the BDLO into a parent DLO and one or more children DLOs.
-Each DLO is discretized into vertices and represented as elastic rods.
-This setup allows DEFT to capture the geometric and physical properties required for dynamic simulation.
-To improve computational efficiency, DEFT then predicts the dynamics of each branch in parallel (Section IV-F).
-During this process, analytical gradients are provided to solve (10), as detailed in Section IV-C, ensuring efficient and stable convergence.
-Next, to address numerical errors, DEFT employs a GNN designed to learn the BDLO’s residual dynamics (Section IV-D).
-By modeling discrepancies between simulated and observed behavior, the GNN refines predictions and enhances overall accuracy.
-After integration, DEFT enforces constraints (Section IV-E) to enforce physical realism.
+Appendix B Supplementary Material for the DEFT Algorithm
+B.1 DEFT Contribution Visualization
+Figure 6: Overview of DEFT contributions.
+First, DEFT partitions the BDLO into parent and child DLOs, each discretized as elastic rods for accurate dynamic simulation (§4.1).
+It then predicts each branch’s dynamics in parallel (§4.5)—using analytical gradients to solve the inner loop for fast, stable convergence (§4.5).
+A GNN learns residual errors between simulated and observed behavior (§4.3), and constraint modules enforce inextensibility, junction attachments, and edge orientations for physical realism (§4.4).
 Inextensibility constraints are applied to each branch, while junction-level constraints ensure proper attachment at branch junctions.
 Additionally, edge orientation constraints enable the propagation of dynamics across these junctions.
-Throughout the entire pipeline, all components remain fully differentiable, allowing for efficient parameter learning from real-world data.
-IV Methodology
-Our objective is to model the dynamic behavior of a BDLO under robotic control.
-Traditional modeling approaches simplify the problem by assuming that the DLO’s endpoints are each held by separate robot end-effectors [8, 28, 16, 10, 29, 30].
-Although this assumption reduces complexity, it limits the potential for dexterous manipulation (e.g., inserting the middle of a BDLO into a plug).
-This work relaxes this assumption and considers scenarios where the BDLO is grasped at an intermediate point, rather than exclusively at its ends.
-This section introduces the key components of BDLO modeling and overviews the various components of DEFT, which are illustrated in Figure 3.
-The key components of DEFT are: 1) analytical gradients to accelerate the optimization procedure used to solve (5); 2) a GNN architecture tailored for residual learning in BDLO simulations; 3) an inextensible constraint formulation that enables the propagation of dynamics at the junction between branches while conserving momentum, which improves the numerical stability of the simulation; and 4) a parallel programming strategy that ensures efficient simulation and gradient computation during training and simulation.
-Please note that to simplify the presentation, we omit the time subscript $t$ in Sections IV-C and IV-E.
-IV-A Branched Deformable Linear Object Model
-Figure 4: An illustration of a branched deformable linear object (BDLO) model. Left: An example configuration of BDLO manipulation. Right: Relevant notation and structure at a junction.
-To model a BDLO, we decompose its structure into a hierarchical arrangement of DLOs.
-The parent DLO is defined as the minimum-length path through the BDLO that connects the two end-effectors (i.e., the points where the structure is manipulated or anchored).
-All other branches diverging from this primary path are modeled as children DLOs that maintain their physical connection to the parent DLO at their attachment points, which we refer to as junction points.
-The parent and children DLOs are each modeled as DLOs and inherit all previously introduced notation.
-For conciseness, we describe the junction point using only three edges at each time instance: two from the parent DLO and one from a child DLO.
-In principle, the methods in DEFT can be extended to handle junctions with more than three edges.
-We distinguish between the parent and child DLOs using superscripts $p$ and $c$, respectively.
-For instance, if a junction is located at the $i$-th vertex, the associated edges are denoted by $\mathbf{e}^{p,i-1}_{t}$, $\mathbf{e}^{p,i}_{t}$, and $\mathbf{e}^{c,1}_{t}$.
-An illustration of a BDLO can be found in Figure 4.
-To capture the influence of bending and twisting, we introduce the moment of inertia of the $i$-th edge which is denoted by $\mathbf{I}^{\text{i}}\in\mathbb{R}^{3\times 3}$.
-It represents how the segment’s mass distribution affects its resistance to rotational deformations.
-The orientation of $\mathbf{e}^{i}_{t}$ is denoted by an angle axis vector $\mathbf{\Omega}^{i}_{t}\in\mathbb{R}^{3}$.
-Let the orientation and position of the two end effectors at time $t$ be denoted by $\mathbf{u}_{t}\in\mathbb{R}^{12}$.
-We refer to $\mathbf{u}_{t}$ as the input.
-We assume that the initial state of the BDLO is fully observable, and that it can be accurately estimated using image segmentation [31] in combination with Gaussian Mixture Models [32].
-IV-B DEFT Overview
-At each time $t$, DEFT simulates the behavior of a BDLO when given predictions of the vertices and their velocities under a specified input as described in Algorithm 1.
-It begins by first associating the grasped vertices with the robot end effector inputs $\mathbf{u}_{t}$.
-DEFT then processes the parent and children DLOs in parallel to compute their respective material frames using (3) (4), which are subsequently used to calculate each branch’s potential energy.
-DEFT then minimizes this potential energy using analytical gradients to determine the optimal parameters $\bm{\theta}_{t}^{*}$ and updates the potential energy accordingly (Line 2).
-These forces are integrated forward in time using a Semi-Implicit Euler method combined with residual learning to update the vertex velocities and positions (Lines 3 and 4).
-However, due to numerical integration errors, branch lengths may change in length.
-In addition, up to this point, DEFT treats each branch’s internal dynamics in isolation, without accounting for junction interactions.
-This omission can hinder the propagation of dynamics across junctions.
-For example, children DLOs may detach from their parent branches, or deformation effects (e.g., bending and twisting) may not propagate through the junction.
-To address these issues, DEFT enforces additional constraints to ensure each branch remains inextensible, maintain branch connections, and allow deformation effects to propagate across the junction (Line 5).
-By applying Algorithm 1 over multiple time increments, DEFT can accurately predict the BDLO’s state over extended time horizons.
-An algorithm overview is illustrated in Figure 3.
-Algorithm 1 $(\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}})=\textbf{DEFT}(\hat{\mathbf{X}}_{\text{t}},\hat{\mathbf{V}}_{\text{t}},\textbf{u}_{t}$)
-1:Associate grasped vertices with $\textbf{u}_{t}$
-2:Calculate Material Frames $\Gamma(\textbf{X}_{t},\bm{\theta}_{t})$triangleright (3),(4)
-3:$\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})=\operatorname*{arg\,min}_{\bm{\theta}_{t}}P(\Gamma(\textbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$triangleright Section.IV-C
-4:$\hat{\textbf{V}}_{t+1}=\hat{\textbf{V}}_{t}-\Delta_{t}\mathbf{M}^{-1}\frac{\partial P(\Gamma(\textbf{X}_{t},\bm{\theta}^{*}_{t}(\textbf{X}_{t},\bm{\alpha})),\bm{\alpha})}{\partial\textbf{X}_{t}}$triangleright (7)
-5:$\hat{\mathbf{X}}_{t+1}=\hat{\mathbf{X}}_{t}+\Delta_{t}\hat{\textbf{V}}_{t+1}+\text{GNN}$triangleright Section.IV-D
-6:Constraints enforcement on $\hat{\mathbf{X}}_{t+1}$triangleright Algorithm 2
-7:$\hat{\textbf{V}}_{t+1}=(\hat{\mathbf{X}}_{t+1}-\hat{\mathbf{X}}_{t})/\Delta_{t}$triangleright Velocity Update
-return $\hat{\mathbf{X}}_{\text{t+1}},\hat{\mathbf{V}}_{\text{t+1}}$
-IV-C Analytical Gradient Derivation
-The goal of this subsection is to present the analytical gradient of $P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})$ with respect to $\bm{\theta}_{t}$.
-To improve the paper’s readability, we summarize the result in the following theorem whose proof can be found in Appendix -A.
-Theorem 1.
-Let $\Gamma(\mathbf{X}_{t},\bm{\theta}_{t})$ be the Material Frame at time $t$, and let $\bm{\alpha}$ contain the bending stiffness $\mathbf{B}\in\mathbb{R}^{3\times 3}$ and twisting stiffness $\beta\in\mathbb{R}$.
-Suppose the total potential energy is the sum of bending and twisting potential energies:
-$\begin{split}P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha})\;=\;P_{\text{bend}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\mathbf{B}\bigr{)}+\\ P_{\text{twist}}\bigl{(}\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\beta\bigr{)}.\end{split}$ (12)
-Then the partial derivative of the bending and twisting potential energies with respect to $\theta^{i}$ are given by
-$\begin{split}\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(&\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\\ &\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},\end{split}$ (13)
-$\begin{split}\frac{\partial P_{\mathrm{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\beta^{i}\,\bigl{(}\theta^{i}-\theta^{i-1}\bigr{)}\;-\;\\ \beta^{i+1}\,\bigl{(}\theta^{i+1}-\theta^{i}\bigr{)},\end{split}$ (14)
-respectively, where the material curvature $\bm{\omega}^{(i,j)}$ is defined as
-$\bm{\omega}^{(i,j)}\;=\;\bigl{(}\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{1}^{j},\;\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{2}^{j}\bigr{)}^{T}\quad\text{for}\quad j\in\{i-1,\,i\},$ (15)
-$\overline{\bm{\omega}}^{(i,j)}$ denotes the undeformed material curvature, and the curvature vector $\kappa\mathbf{b}^{i}$ in (15) is
-$\kappa\,\mathbf{b}^{i}\;=\;\frac{2\,\bigl{(}\mathbf{e}^{i-1}\times\mathbf{e}^{i}\bigr{)}}{\|\mathbf{e}^{i-1}\|_{2}\,\|\mathbf{e}^{i}\|_{2}\;+\;\mathbf{e}^{i-1}\cdot\mathbf{e}^{i}}.$ (16)
-IV-D Integration Method with Residual Learning
-This subsection constructs the GNN for residual learning in (11).
-The motivation for employing a GNN lies in its ability to represent a BDLO as a graph, where each vertex is treated as a node and edges capture both intra-branch connections and junctions.
-This graph-based formulation translates into an adjacency matrix $\mathbf{A}\in\mathbb{R}^{(n^{p}+n^{c})\times(n^{p}+n^{c})}$, which links individual branch nodes and explicitly encodes node information across branch junctions.
-This adjacency matrix is defined as:
-$\mathbf{A}_{i_{1},i_{2}}=\begin{cases}1&\text{if there is an edge between }i_{1}\text{ and }i_{2},\\ 0&\text{otherwise,}\end{cases}$ (17)
-An illustration of $\mathbf{A}$ can be found in Figure 5.
-In practice, DEFT constructs $\mathbf{A}$ for each BDLO whose behavior a user wants to predict.
-This adjacency matrix is utilized within a Graph Convolutional Network (GCN) [33] to aggregate learned features.
-A single GCN layer is defined as
-$\text{GCN}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})=\mathbf{\tilde{A}}\,\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\,\mathbf{W},$ (18)
-where $\mathbf{\tilde{A}}$ is the normalized adjacency matrix
-$\mathbf{\tilde{A}}=\mathbf{D}^{-\frac{1}{2}}\mathbf{A}\mathbf{D}^{-\frac{1}{2}}.$ (19)
-Note that $\mathbf{\tilde{A}}$ is the normalization of $\mathbf{A}$ using the degree matrix $\mathbf{D}=\mathrm{diag}(d_{1},d_{2},\ldots,d_{n})$ where $d_{i1}=\sum_{i2}A_{i1,i2}$.
-This normalization ensures that nodes with a high degree do not disproprtionately influence the overall performance.
-This ensures consistent feature magnitudes across the learned representation [34].
-The node feature matrix, $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})\in\mathbb{R}^{(n^{p}+n^{c})\times h}$, is constructed by concatenating its arguments together.
-The trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ is implemented via a multi-layer perceptron (MLP) [35], which projects input features into latent representations.
-Figure 5: The adjacency matrix $\mathbf{A}$ for the junction in Figure 4 captures both self-loops (diagonal entries) and inter-node connections (off-diagonal entries).
-By embedding each node’s local dynamics alongside its coupling to neighboring nodes, $\mathbf{A}$ enables an graph representation of the BDLO’s behavior.
-IV-E Enforcing Constraints
-After performing numerical integration, DEFT applies constraints to ensure branch inextensibility, ensure that branches are attached at the junctions, and ensure that deformations are propagated throughout the BDLO.
-To accomplish this, DEFT updates the positions of each vertex using correction terms $\Delta\hat{\mathbf{X}}_{\text{}}$.
-One could just update the positions of each vertex by just projecting directly onto the constraint set.
-However, this induces highly nonlinear changes in momentum and can lead to simulation instability and unrealistic behavior.
-Motivated by the PDB method [36], DEFT enforces the aforementioned constraints while trying to preserve momentum.
-DEFT does this by formulating the computation of the correction term as the solution to an optimization problem.
-As we illustrate through our experimental evaluation in Section V, this leads to more stable simulations, which corresponds to more realistic predictions.
-Finally, note that DEFT could formulate the optimization problem to compute the correction terms as a single large optimization problem that considers constraints between all vertices in the BDLO simultaneously.
-However, this large optimization problem can be difficult to solve rapidly.
-Instead DEFT enforces these constraints sequentially between neighboring nodes.
-Next, we summarize each of these optimization problems.
-IV-E1 Inextensibility Enforcement
+All components remain fully differentiable, enabling efficient parameter learning from real-world data.
+B.2 Definition of $\hat{\bm{\Omega}}^{i}_{\Delta,t}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ and $\hat{\bm{\Omega}}^{i}$
+In practice, it can be both challenging and impractical to attach hardware onto each segment for tracking its orientation changes.
+To address this limitation, we approximate the change in orientation, as illustrated in Figure 7.
+We define $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ as follows:
+$\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\frac{\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}}{||\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}||_{2}}\times\frac{(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})-(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})}{||(\hat{\mathbf{x}}^{i+1}+\Delta\hat{\mathbf{x}}^{i+1})-(\hat{\mathbf{x}}^{i}+\Delta\hat{\mathbf{x}}^{i})||_{2}}$ (18)
+Intuitively, the cross product of each edge’s tangent vector provides an approximation of the rotation axis and orientation change for the segment $i$.
+This approximation becomes more accurate when $\Delta_{t}$ is small.
+Next, we describe how to compute $\hat{\bm{\Omega}}^{i}$.
+Note that it is defined recursively using the computation at the previous time step.
+As a result just within this appendix, we add a subscript $t$ to the symbol.
+To compute $\hat{\bm{\Omega}}^{i}_{t+1}$, we first convert both $\hat{\bm{\Omega}}^{i}_{t}$ and the correction rotation $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}}_{t},\Delta\hat{\mathbf{X}}_{t})$ from angle-axis to quaternions.
+We then update $\hat{\bm{\Omega}}^{i}_{t}$ by applying $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}}_{t},\Delta\hat{\mathbf{X}}_{t})$ in quaternion form, and finally convert the result back to angle-axis coordinates to obtain $\hat{\bm{\Omega}}^{i}_{t+1}$.
+Note that we do not continuously use quaternions because angle–axis is more convenient for representing angular momentum.
+Figure 7: An illustration of $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$
+B.3 Residual Learning Details
+The node feature matrix $\mathbf{F}_{t}(\hat{\mathbf{X}}_{t},\hat{\mathbf{V}}_{t},\bm{\alpha})$ is constructed by concatenating its arguments together.
+The trainable weight matrix $\mathbf{W}\in\mathbb{R}^{h\times 3}$ is implemented using a multi-layer perceptron (MLP) [38], consisting of two linear layers with a ReLU activation in between and a feature dimension of 64 for message passing.
+This projects input features into latent representations.
+B.4 Supplementary Material for Enforcing Constraints
+In what follows, we define $C(\hat{\mathbf{z}})$ for each constraint in Theorem 1, present the solutions as individual theorems with proofs, and summarize their enforcement in Algorithm 2.
+B.4.1 Inextensibility Enforcement
 To enforce inextensibility of each branch, we define a constraint function between successive vertices.
 This function penalizes deviations of each segment’s length from its undeformed length.
-By applying the following theorem, whose proof can be found in [8, Appendix B.2], one can compute $\Delta\hat{\mathbf{x}}^{i},\Delta\hat{\mathbf{x}}^{i+1}$ explicitly:
+By applying the following theorem, one can compute $\Delta\hat{\mathbf{x}}^{i},\Delta\hat{\mathbf{x}}^{i+1}$ explicitly:
 Theorem 3 (Computing the Corrections to Enforce Inextensibility).
@@
 The correction $\Delta\hat{\mathbf{X}}$ to enforce inextensibility at joint $i$ can be computed by solving the following optimization problem:
-$\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (21)
-s.t. $\hskip 8.5359pt\Delta\hat{\mathbf{x}}^{j}=0$ $\forall j\notin\{i,i+1\},$
-$\hskip 8.5359pt\mathbf{M}^{i}\Delta\hat{\mathbf{x}}^{i}+\mathbf{M}^{i+1}\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{0},$
-$\hskip 8.5359pt\mathbf{I}^{i}\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0},$
-where $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\in\mathbb{R}^{3}$ is defined in Appendix -B and represents the orientation change of the $i$th edge arising from $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ and $\mathbf{0}$ correspond to the zero vector of appropriate size.
+$\underset{\Delta\hat{\mathbf{X}}}{\min}\quad$ $\frac{1}{2}\left(C^{i}_{I}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (20)
+s.t. $\Delta\hat{\mathbf{x}}^{j}=0,\quad\forall j\notin\{i,i+1\},$
+$\mathbf{M}^{i}\Delta\hat{\mathbf{x}}^{i}+\mathbf{M}^{i+1}\Delta\hat{\mathbf{x}}^{i+1}=\mathbf{0},$
+$\mathbf{I}^{i}\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0}.$
+where $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\in\mathbb{R}^{3}$ is defined in Appendix B.2 and represents the orientation change of the $i$th edge arising from $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ and $\mathbf{0}$ correspond to the zero vector of appropriate size.
 The only non-zero elements of the minimizer to this optimization problem can be computed explicitly and are equal to:
@@
 For convenience, let the function that computes the minimizer of optimization problem (20) using $\hat{\mathbf{X}}$ be denoted by $D^{i}_{I}$ (i.e., $D^{i}_{I}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
-IV-E2 Enforcing Attachment at the Junction
+Proof. The proof can be found in [14, Appendix B.2].
+B.4.2 Enforcing Attachment at the Junction
 To ensure the child branch remains attached to the parent branch, we define a constraint function between their junction vertices.
-By applying the following theorem, whose proof can be found in Appendix -C, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
+By applying the following theorem, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
 Theorem 4.
@@
 The only non-zero elements of the minimzer to this optimization problem can be computed explicitly and are equal to:
-$\Delta\hat{\mathbf{x}}^{p,i}=\mathbf{M}^{c,1}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}},$
-$\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{M}^{p,i}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}\ C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{p,i}-\hat{\mathbf{x}}^{c,1})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}}.$
+$\Delta\hat{\mathbf{x}}^{p,i}=\mathbf{M}^{c,1}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}},$ (24)
+$\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{M}^{p,i}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}\ C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{p,i}-\hat{\mathbf{x}}^{c,1})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}}.$ (25)
 Note that the junction attachment constraint is satisfied at vertex $p,i$ when $C^{p,i}_{I}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$.
 For convenience, let the function that computes the minimizer of optimization problem (20) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{A}$ (i.e., $D^{p,i}_{A}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
-IV-E3 Edge Orientation Constraints at the Junction
+Proof. To prove Theorem 4, we first formulate the optimization problem:
+$\underset{\Delta\hat{\mathbf{X}}}{\min}\quad$ $\frac{1}{2}\left(C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (26)
+s.t. $\Delta\hat{\mathbf{x}}^{j}=\mathbf{0},\quad\forall j\notin\{(p,i),(c,1)\},$ (27)
+$\mathbf{M}^{p,i}\,\Delta\hat{\mathbf{x}}^{p,i}+\mathbf{M}^{c,1}\,\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{0},$ (28)
+$\mathbf{I}^{p,i}\,\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\mathbf{0}.$ (29)
+To solve the above optimization, we introduce Lagrange multipliers $\bm{\lambda}_{\text{l}}\in\mathbb{R}^{3}$ and $\bm{\lambda}_{\text{r}}\in\mathbb{R}^{3}$ associated with constraints (28) and (29), respectively.
+The corresponding Lagrangian $\mathcal{L}$ can be found as follow:
+$\begin{split}\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta&\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})=\frac{1}{2}\left(C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}-\\ &\bm{\lambda}^{T}_{\text{l}}(\mathbf{M}^{i}\Delta\hat{\textbf{x}}^{p,i}+\mathbf{M}^{c,1}\Delta\hat{\mathbf{x}}^{c,1})-\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\end{split}$ (30)
+Next, we take the partial derivatives of (30) with respect to $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}}$.
+Setting each derivative to zero yields the system of equations (31), (32), (33), and (34).
+$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\Delta\hat{\mathbf{x}}^{p,i}}=\\ &C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\frac{(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})}{||(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})||_{2}}-\bm{\lambda}^{T}_{\text{l}}\mathbf{M}^{p,i}-\frac{\partial\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})}{\partial\Delta\hat{\mathbf{x}}^{p,i}}=\textbf{0}\end{split}$ (31)
+$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\Delta\hat{\mathbf{x}}^{c,1}}=\\ &C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\frac{(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})-(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})}{||(\hat{\textbf{x}}^{p,i}+\Delta\hat{\textbf{x}}^{p,i})-(\hat{\mathbf{x}}^{c,1}+\Delta\hat{\mathbf{x}}^{c,1})||_{2}}-\bm{\lambda}^{T}_{\text{l}}\mathbf{M}^{c,1}-\frac{\partial\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})}{\partial\Delta\hat{\mathbf{x}}^{c,1}}=\textbf{0}\end{split}$ (32)
+$\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\bm{\lambda}_{\text{l}}}=-(\mathbf{M}^{i}\Delta\hat{\textbf{x}}^{p,i}+\mathbf{M}^{c,1}\Delta\hat{\mathbf{x}}^{c,1})=\textbf{0}$ (33)
+$\begin{split}&\frac{\partial\mathcal{L}(\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1},\bm{\lambda}_{\text{l}},\bm{\lambda}_{\text{r}})}{\partial\bm{\lambda}_{\text{r}}}=-\textit{{I}}^{i}\bm{\lambda}^{T}_{\text{r}}\mathbf{I}^{p,i}\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}\end{split}$ (34)
+We first observe that solving (33) leads to the following equation:
+$\Delta\hat{\textbf{x}}^{c,1}=-(\mathbf{M}^{c,1})^{-1}\mathbf{M}^{p,i}\Delta\hat{\mathbf{x}}^{p,i}$ (35)
+This indicates that $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ are colinear.
+Based on the definition of $\hat{\bm{\Omega}}^{i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})$ in Appendix B.2, one can see that if $\Delta\hat{\textbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ are colinear, then $\hat{\bm{\Omega}}^{p,i}_{\Delta}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}$.
+Note that this is true regardless of the size of $\Delta_{t}$.
+Hence, it naturally satisfies (34).
+Next, summing (31) with (32) yields the following:
+$\bm{\lambda}_{\text{l}}\mathbf{M}^{p,i}+\bm{\lambda}_{\text{l}}\mathbf{M}^{c,1}=\textbf{0}$ (36)
+As mass matrices are always positive definite.
+This indicates $\bm{\lambda_{l}}=\textbf{0}$.
+Based on (31), (35) and (36), solving for $\Delta\hat{\mathbf{x}}^{p,i}$ is equivalent to solving the following equation:
+$C^{p,i}_{A}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\textbf{0}$ (37)
+After applying some algebraic manipulations, $\Delta\hat{\mathbf{x}}^{p,i}$ can be found to be:
+$\Delta\hat{\mathbf{x}}^{p,i}=\mathbf{M}^{c,1}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}},$
+Using $\Delta\hat{\mathbf{x}}^{p,i}$ and (35), one can find:
+$\Delta\hat{\mathbf{x}}^{c,1}=\mathbf{M}^{p,i}(\mathbf{M}^{p,i}+\mathbf{M}^{c,1})^{-1}\ C^{p,i}_{A}(\hat{\mathbf{X}},\mathbf{0})\frac{(\hat{\mathbf{x}}^{p,i}-\hat{\mathbf{x}}^{c,1})}{||\hat{\mathbf{x}}^{c,1}-\hat{\mathbf{x}}^{p,i}||_{2}}.$
+B.4.3 Edge Orientation Constraints
 Many practical wire-harnesses feature junctions that are nearly rigid.
@@
 To capture this behavior, we model each edge as a rigid body and define a constraint function that computes the relative orientation of each body.
-We then require that this constraint function is approximately satisfied where ap By applying the following theorem, whose proof can be found in Appendix -D, one can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
+We then require that this constraint function is approximately satisfied.
+One can compute $\Delta\hat{\mathbf{x}}^{p,i},\Delta\hat{\mathbf{x}}^{c,1}$ explicitly:
 Theorem 5.
@@
 Let the constraint function to enforce the orientation of the junction be defined as:
-$C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\big{)}-\\ \big{(}\hat{\bm{\Omega}}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}}\big{)}\|_{2}-\epsilon,$ (25)
-where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix -B and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid.
+$C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})=\|\big{(}\hat{\bm{\Omega}}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{p,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\big{)}-\big{(}\hat{\bm{\Omega}}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})+\hat{\bm{\Omega}}_{\Delta}^{c,i}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}}\big{)}\|_{2}-\epsilon,$ (38)
+where $\hat{\bm{\Omega}}_{\Delta}$ and $\hat{\bm{\Omega}}$ are defined in Appendix B.2 and $\epsilon>0$ bounds the mismatch between the updated parent–child orientation and the original orientation, ensuring that each junction remains almost rigid.
 The orientation correction at the junction is obtained by solving the following optimization problem:
 $\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (39)
-s.t. $\hskip 8.5359pt\mathbf{I}^{p,i}\hat{\bm{\Omega}}_{\Delta}^{p,i}+\mathbf{I}^{c,1}\hat{\bm{\Omega}}_{\Delta}^{c,1}=\mathbf{0}$
-When $\Delta_{t}$ is sufficiently small, then one can compute explicit formulas for $\Delta\hat{\mathbf{X}}$ which can be found in Appendix -D.
+s.t. $\hskip 8.5359pt\mathbf{I}^{p,i}\hat{\bm{\Omega}}_{\Delta}^{p,i}+\mathbf{I}^{c,1}\hat{\bm{\Omega}}_{\Delta}^{c,1}=\mathbf{0}$ (40)
+When $\Delta_{t}$ is sufficiently small, explicit formulas for $\Delta\hat{\mathbf{X}}$ can be computed as follows:
+$\hat{\bm{\Omega}}_{\Delta}^{p,i}\;=\;\mathbf{I}^{c,1}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}},$ (41)
+$\hat{\bm{\Omega}}_{\Delta}^{c,1}\;=\;\mathbf{I}^{p,i}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{p,i}-\hat{\bm{\Omega}}^{c,1}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}}.$ (42)
+Once $\hat{\bm{\Omega}}_{\Delta}^{p,i},\hat{\bm{\Omega}}_{\Delta}^{c,1}$ are obtained, an approach described in the following are used to update $\hat{\mathbf{x}}^{p,i}$ and $\hat{\mathbf{x}}^{c,1}$ accordingly.
 Note that the junction orientation constraint is satisfied at vertex $p,i$ when $C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})^{2}=0$.
 For convenience, let the function that computes the minimizer of optimization problem (20) using $\hat{\mathbf{X}}$ be denoted by $D^{p,i}_{O}$ (i.e., $D^{p,i}_{O}:\hat{\mathbf{X}}\mapsto\Delta\hat{\mathbf{X}}$).
-IV-E4 Algorithm to Implement Corrections Due To Constraint Enforcement
+Proof.
+Figure 8: An illustration of applying $\hat{\bm{\Omega}}_{\Delta}^{i}$ to rotate the $i$-th segment.
+For convenience in the proof, we suppress the arguments to functions for convenience.
+To prove Theorem 5, we begin by describing how $\hat{\bm{\Omega}}^{i}_{\Delta}$ is applied to rotate the $i$-th segment.
+Note that $\Delta_{t}$ should be sufficiently small for $\hat{\bm{\Omega}}^{i}_{\Delta}$ to accurately approximate the change in orientation.
+As shown in Figure 8, the rotation axis coincides with the center of the $i$-th segment.
+In practice, the rotation of the $i$-th segment is achieved by converting $\hat{\bm{\Omega}}^{i}_{\Delta}$ into the rotation matrix $\hat{\bm{\Omega}}^{i}_{\Delta}\mapsto\mathbf{R}^{i}_{\Delta}$.
+Consequently, we obtain $\Delta\hat{\mathbf{x}}^{i}$ and $\Delta\hat{\mathbf{x}}^{i+1}$ via
+$\Delta\hat{\mathbf{x}}^{i}\;=\;\mathbf{R}^{i}_{\Delta}\,\frac{\bigl{(}\hat{\mathbf{x}}^{i+1}-\hat{\mathbf{x}}^{i}\bigr{)}}{2},$ (43)
+$\Delta\hat{\mathbf{x}}^{i+1}\;=\;\mathbf{R}^{i}_{\Delta}\,\frac{\bigl{(}\hat{\mathbf{x}}^{i}-\hat{\mathbf{x}}^{i+1}\bigr{)}}{2}.$ (44)
+Because each segment rotates only about its own center, linear momentum is naturally conserved. Thus, we can simplify the optimization problem and write the following optimization problem:
+$\underset{\Delta\hat{\mathbf{X}}}{\min}$ $\hskip 8.5359pt\frac{1}{2}\left(C^{p,i}_{O}(\hat{\mathbf{X}},\Delta\hat{\mathbf{X}})\right)^{2}$ (45)
+s.t. $\hskip 8.5359pt\mathbf{I}^{p,i}\hat{\bm{\Omega}}_{\Delta}^{p,i}+\mathbf{I}^{c,1}\hat{\bm{\Omega}}_{\Delta}^{c,1}=\mathbf{0}$
+where (38), (39), and (40) are analogous to (23), (26), and (28), respectively.
+Next, we introduce a Lagrange multiplier and follow the same derivation steps as in (31), (35), and (36), which leads to:
+$\hat{\bm{\Omega}}_{\Delta}^{p,i}\;=\;\mathbf{I}^{c,1}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}},$
+$\hat{\bm{\Omega}}_{\Delta}^{c,1}\;=\;\mathbf{I}^{p,i}\,\bigl{(}\mathbf{I}^{c,1}+\mathbf{I}^{p,i}\bigr{)}^{-1}\,C^{p,i}_{O}(\hat{\mathbf{X}},\mathbf{0})\,\frac{\hat{\bm{\Omega}}^{p,i}-\hat{\bm{\Omega}}^{c,1}}{\|\hat{\bm{\Omega}}^{c,1}-\hat{\bm{\Omega}}^{p,i}\|_{2}}.$
+Once $\hat{\bm{\Omega}}_{\Delta}^{p,i},\hat{\bm{\Omega}}_{\Delta}^{c,1}$ are obtained, (43) and (44) are used to update $\hat{\mathbf{x}}^{p,i}$ and $\hat{\mathbf{x}}^{c,1}$ accordingly.
+B.4.4 Algorithm to Implement Corrections Due To Constraint Enforcement
 As described earlier, we run each of these optimization problems sequentially to generate a correction to the output of the integration method (Line 5 in Algorithm 1).
@@
 11:return : $\hat{\textbf{X}}_{t}$triangleright Updated Vertices
-IV-F Parallel Programming
-We implement DEFT using PyTorch [26], taking advantage of its batch operations to execute Algorithm 1.
-PyTorch’s support for batch-wise computations enables parallel processing of multiple parent and child branch instances, significantly enhancing computational efficiency as we illustrate empirically in Section V-D.
-Specifically, the first batch is dedicated to executing the parent branch’s algorithm, while the remaining batches handle executions for the children branches.
-One challenge arises when the number of nodes in the parent branch ($n^{p}$) differs from the number in a child branch ($n^{c}$).
-This dimensional mismatch prevents the direct construction of uniform matrices for batch operations.
-To overcome this, we assume $n^{p}>n^{c}$ and introduce padded matrices and vertices by appending zeros to the child branch’s structure to match the parent branch’s dimensions.
-This strategy ensures uniform matrix construction across the batches, which facilitates seamless batch-wise operations.
-During parameter updates, we mask out the gradients from padded elements to prevent them from influencing the learned parameters.
-V Experiments and Results
-This section presents the real-world evaluation of DEFT, covering: 1) A quantitative accuracy comparison with state-of-the-art methods, including scenarios where grasping occurs at various points along the BDLO, such as parent-branch ends and midpoints; 2) An assessment of DEFT’s inference speed; 3) An ablation study exploring parameter auto-tuning, residual learning, and a naive multi-branch approach that connects multiple single-DLO models; 4) Shape-matching manipulation demonstrations, highlighting DEFT’s real-world manipulation efficacy; 5) Thread insertion manipulation demonstrations, showcasing DEFT’s versatility.
-Code and data can be found at https://github.com/roahmlab/DEFT.
-To the best of our knowledge, this will be the first publicly available dataset and code for modeling BDLOs.
-V-A Hardware Setup
-Figure 6: An illustration of experiment setup.
-Top: Data collection setup with a dual-arm robot.
-Bottom: An illustration of the BDLOs used to evaluate and compare the performance of DEFT against baseline methods.
-To increase dataset diversity, we vary the BDLOs’ stiffness, length, weights, number of vertices, and junction structures.
-TABLE I: A summary of each BDLO’s material properties and marker count used in the real-world experiment. Stiffness is estimated based on a relative scale.
-BDLO Parent Stiffness / #Markers Child1 Stiffness / #Markers Child2 Stiffness / #Markers Dataset 1 Time (s) Dataset 2 Time (s)
-1 Medium / 13 Medium / 4 Medium / 3 505 440
-2 Low / 12 Low / 4 Low / 4 735 N/A
-3 Medium / 12 Low / 2 High / 3 645 445
-4 Medium / 12 High / 2 High / 2 495 N/A
-V-A1 Hardware Setup
+B.5 Proof of Theorem 2
+To prove the theorem, we begin by introducing the curvature binormal and material curvature.
+Material Curvatures: The curvature binormal $\kappa\mathbf{b}^{i}$ is traditionally used to represent the turning angle and axis between two consecutive edges:
+$\kappa\mathbf{b}^{i}=\frac{2\,\mathbf{e}^{i-1}\times\mathbf{e}^{i}}{\|\mathbf{e}^{i-1}\|_{2}\|\mathbf{e}^{i}\|_{2}+\mathbf{e}^{i-1}\cdot\mathbf{e}^{i}},$ (46)
+where $\mathbf{e}^{i-1}$ and $\mathbf{e}^{i}$ are consecutive edge vectors, $\times$ denotes the cross product, and $\cdot$ denotes the dot product.
+By incorporating the material frame (12) and (13), the curvature binormal (46) is projected onto the material frame to quantify the extent to which the curvature aligns with the frame’s orientation.
+This projection, which we call the material curvature, provides additional insights into the deformation characteristics, allowing us to distinguish between bending and twisting behaviors within the material frame and is defined as:
+$\bm{\omega}^{(i,j)}=\left(\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{1}^{j},\kappa\mathbf{b}^{i}\cdot\mathbf{m}_{2}^{j}\right)^{T}\quad\text{for}\quad j\in\{i-1,i\}.$ (47)
+Potential Energy: The potential energy is composed of the bending energy and twisting energy:
+$P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=P_{\text{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}+P_{\text{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)},$ (48)
+where
+$\begin{split}&P_{\text{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=\\ &\sum_{i=1}^{n-1}\sum_{j=i-1}^{i}\frac{1}{2}\left(\bm{\omega}^{(i,j)}-\overline{\bm{\omega}}^{(i,j)}\right)^{T}\mathbf{B}^{j}\left(\bm{\omega}^{(i,j)}-\overline{\bm{\omega}}^{(i,j)}\right),\end{split}$ (49)
+and
+$P_{\text{twist}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}=\sum_{i=1}^{n-1}\frac{1}{2}\beta^{i}\left(\theta^{i}-\theta^{i-1}\right)^{2},$ (50)
+where $\overline{\bm{\omega}}^{(i,j)}$ denotes the undeformed material curvature, which is calculated when the DLO is in a static state without any external or internal forces applied, and $\mathbf{B}$ and $\beta$, are components of $\bm{\alpha}$, representing the bending stiffness and twisting stiffness, respectively.
+Gradient of Potential Energy: With (49) and (50), the gradient of (48) can be derived as following:
+$\frac{\partial P(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}+\frac{\partial P_{twist}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}$ (51)
+We begin by deriving the first term. Note that $\theta^{i}$ is only relevant in $\bm{\omega}^{(i,i)}$ and $\bm{\omega}^{(i+1,i)}$. Using the chain rule, we obtain:
+$\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\sum_{k=i}^{i+1}\frac{\partial P_{bend}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\bm{\omega}^{(k,i)}}\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}.$ (52)
+To compute $\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}$, we use the identities $\frac{\partial\mathbf{m}_{1}^{i}}{\partial\theta^{i}}=\mathbf{m}_{2}^{i}$ and $\frac{\partial\mathbf{m}_{2}^{i}}{\partial\theta^{i}}=-\mathbf{m}_{1}^{i}$.
+These lead to:
+$\frac{\partial\bm{\omega}^{(k,i)}}{\partial\theta^{i}}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\bm{\omega}^{(k,i)}$ (53)
+Substituting this result back, we obtain:
+$\frac{\partial P_{\mathrm{bend}}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}\;=\;\sum_{k=i}^{\,i+1}\bigl{(}\mathbf{B}^{k}\,(\bm{\omega}^{(k,i)}-\bar{\bm{\omega}}^{(k,i)})\bigr{)}^{T}\cdot\begin{bmatrix}0&1\\[6.0pt] -1&0\end{bmatrix}\bm{\omega}^{(k,i)},$ (54)
+The second term of (51) can be derived, resulting in:
+$\frac{\partial P_{twist}(\Gamma(\mathbf{X}_{t},\bm{\theta}_{t}),\bm{\alpha}\bigr{)}}{\partial\theta^{i}}=\beta^{i}(\theta^{i}-\theta^{i-1})-\beta^{i+1}(\theta^{i+1}-\theta^{i})$ (55)
+By substituting (54) and (55) into (51), we obtain the analytical gradient of the potential energy.
+Appendix C Supplementary Material for Experiments
+C.1 Hardware Setup
 Four distinct BDLOs were constructed to validate DEFT’s modeling accuracy as shown at the bottom of Figure 9.
-The material stiffness description of each BDLO is outlined in Table I.
+The material stiffness description of each BDLO is outlined in Table 4.
 Note that we introduced stiff child branches attached to the parent branch in BDLO3 and BDLO4.
@@
 These attachments help ensure secure mounting, preserve structural integrity, and manage loads in real‐world applications.
+Note that we selected two child branches for each BDLO to evaluate DEFT’s performance across combinations of branches with varying stiffness.
+However, our methodology is theoretically scalable to BDLOs with arbitrary tree-like topologies.
 Spherical markers were attached to each BDLO to facilitate dataset creation for both training and evaluation, with the number of vertices set equal to the number of MoCap markers.
@@
 Note in our implementation of DEFT we set $\epsilon=0.1$ (Theorem 5) and $\kappa=0.02$ (Algorithm 2).
-V-A2 Dataset Collection and Training
-We collected dynamic trajectory data for selected BDLOs in the real world using a motion capture system operating at 100 Hz, capturing both slow and fast BDLO motions, as illustrated at the top of Figure 6.
-We split our data collection into two categories:
-  • Dataset 1: each robot grasps one end of the BDLO. Data from this category were collected for all BDLOs.
-  • Dataset 2: One robot grasped one end of the BDLO while the other robot grasped the BDLO at its midpoint. This dataset is collected for BDLOs 1 and 3.
-Total collection times are provided in Table I.
+The parameters were selected to balance stability, convergence speed, and accuracy.
 The dataset is split with $\sim$75for training and $\sim$25for evaluation.
-The model for each of the four BDLOs is trained separately using the trajectory data collected from our real-world experiments.
-During training, models are trained using the same 100 steps (i.e., 1s) of dynamic motion Each method’s performance is evaluated when predicting the state of the wire over a 500 step (i.e.
-5s) prediction horizon using just the state of the wire at the initial time and the orientation and position of the two end effectors at each time step.
-V-B Baseline Comparisons
-Only a limited number of papers have investigated the dynamic behavior of BDLOs, and none provide a differentiable framework for learning real-world dynamics from real-world datasets.
-Moreover, no public code repositories are available for these approaches, making direct replication infeasible.
-We select two available implementations, Tree-LSTM [38] and GCN [33] as baseline prediction models for BDLOs.
-Tree-LSTM and GCN both capture structured relationships.
-Specifically, Tree-LSTM leverages hierarchical tree representations, while GCN excels in graph-based modeling.
-Because BDLOs can be described by nodes and edges, these architectures are naturally able to incorporate a BDLO’s dynamics.
-V-C Modeling Results
-Table II summarizes the average RMSE loss when comparing DEFT to baseline methods for two categories of dataset.
-Performance is evaluated over a 5 second prediction horizon for four different BDLOs.
-Note that DEFT outperforms all baselines across all four BDLOs.
-An illustration of the performance of DEFT, Tree-LSTM, and an ablation of DEFT, which is described in more detail in the next subsection, can be found in Figure 7.
-A video of related experiments can be found in the supplementary material.
-TABLE II: Modeling Results Over a $5$s Horizon
-Modeling Accuracy (RMSE, $10^{-2}$ m)
-Dataset 1 Dataset 2
-Method 1 2 3 4 1 3
-GCN [33] 6.70 6.42 5.25 4.20 7.33 4.24
-Tree-LSTM [38] 5.32 4.50 2.46 2.71 7.08 4.55
-DEFT 1.87 2.82 1.51 1.41 1.49 1.80
+Figure 9: An illustration of experiment setup.
+Left: Data collection setup with a dual-arm robot.
+Right: An illustration of the BDLOs used to evaluate and compare the performance of DEFT against baseline methods.
+To increase dataset diversity, we vary the BDLOs’ stiffness, length, weights, number of vertices, and junction structures.
+Table 4: A summary of each BDLO’s material properties and marker count used in the real-world experiment. Stiffness is estimated based on a relative scale.
+BDLO Parent Stiffness / #Markers Child1 Stiffness / #Markers Child2 Stiffness / #Markers Dataset 1 Time (s) Dataset 2 Time (s)
+1 Medium / 13 Medium / 4 Medium / 3 505 440
+2 Low / 12 Low / 4 Low / 4 735 N/A
+3 Medium / 12 Low / 2 High / 3 645 445
+4 Medium / 12 High / 2 High / 2 495 N/A
+C.2 Training Setup
+Let $\mathbf{U}_{1:T-1}$ denote the set of inputs applied from time $t=1$ to $t=T-1$, and let $\mathbf{X}_{1:T}=\{\mathbf{X}_{1},\mathbf{X}_{2},\dots,\mathbf{X}_{T}\}\in\mathbb{R}^{T\times n\times 3}$ denote the corresponding ground-truth trajectory of the BDLO from $t=1$ to $t=T$.
+Given known initial conditions $\mathbf{X}_{1}$, $\mathbf{V}_{1}$, and inputs $\mathbf{U}_{1:T-1}$, Algorithm 1 can be recursively applied to predict the trajectory $\hat{\mathbf{X}}_{2:T}$.
+Let $\bm{\phi}$ denote the parameters of the GNN.
+The training objective is formulated as the following optimization problem:
+$\underset{\bm{\alpha},\bm{\phi}}{\min}\sum_{t=1}^{T-1}\|\mathbf{X}_{t+1}-\hat{\mathbf{X}}_{t+1}\|_{2}$ (56)
+Leveraging DEFT’s differentiability, $T$ can be selected as greater than 2 to better capture long-term dynamics.
+This multi-step training approach results in improved prediction accuracy compared to single-step training, as demonstrated in Table 5.
+Table 5: Ablation Study with Dataset 1
+Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
+Dataset 1
+Method           1          2
+Single-step Training           2.22          3.41
+Multi-step Training           1.87           2.82
+C.3 Computational Speed
+This section evaluates the computational efficiency of DEFT when compared to baseline methods.
+It also examines how employing the analytical gradient derived from Theorem 2 to compute (3) and the parallel programming approach detailed in Section 4.5, impacts overall computational time.
+All experiments in this section were conducted in Python on an Ubuntu 20.04 equipped with an AMD Ryzen PRO 5995WX CPU, 256GB of RAM, and 128 cores.
+C.3.1 Computational Speed with Analytical Gradient
+Table 6: Compuational Time for (3) with Dataset 1
+Time ($10^{-2}$ s)
+Method 1 2 3 4
+Analytical Gradient 0.34 0.27 0.41 0.46
+Numerical Gradient 0.74 0.65 0.84 0.91
+We compare the computational speed of using analytical gradients when compared to using numerical gradients.
+As summarized in Table 6, computing (3) with the analytical gradient is approximately twice as fast as using numerical gradients.
+C.3.2 Computational Speed with Parallel Programming
+Figure 10: Computational speed comparison: parallel programming versus a naive (sequential) implementation.
+We compare the computational performance of using parallel programming approach described in Section 4.5 to a naive (sequential) implementation, in which each branch is simulated one after the other.
+In this experiment, we use the material properties of BDLO1 and gradually increase the number of children branches from 1 to 8.
+As shown in Figure 10, the naive implementation’s computation time grows more rapidly than that of the parallel approach.
+However, a certain increase in computational time is still expected with parallelization.
+As illustrated in Figure 6, constraints at each junction between the parent branch and its child branches prevent a fully batch-based operation, thereby resulting in an increase in the computational time.
+C.4 Ablation Study on Residual Learning Architecture
+To further demonstrate the significance of using GNN for residual learning, we replace it with the fully connected MLP described in Appendix B.3 and test on BDLO 2 and 3 from Dataset 1.
+We use BDLO 2 and 3 from dataset 1.
+As shown in Table 7, residual learning using a naive MLP is running slightly more efficient but provides limited improvements when compared to the proposed GNN approach.
+This illustrates the importance of the GNN because even modest improvements are important for prediction accuracy in long-term horizon predictions, where small errors can be significantly amplified during extended open-loop predictions.
+Table 7: Ablation Study Comparing GNN and MLP Residual Learning Approaches
+Accuracy ($10^{-2}$ m) Comp. Time ($10^{-2}$ s)
+Method BDLO 2 BDLO 3 BDLO 2 BDLO 3
+Replace GNN with MLP 3.11 1.64 0.67 0.72
+DEFT 2.82 1.51 0.70 0.76
+C.5 Modeling and Planning Visualization
+Additional visulization of modeling and planning results can be found in Figure 11 and Figure 12 respectively.
 Figure 11: Visualization of the predicted trajectories for BDLO 1 under two manipulation scenarios, using DEFT, a DEFT ablation that leaves out the constraint described in Theorem 5, and Tree-LSTM.
@@
 The prediction is performed recursively, without requiring additional ground-truth data or perception inputs throughout the entire process.
-V-D Compuational Experiment
-TABLE III: One Step Inference Time with Dataset 1
-Time ($10^{-2}$ s)
-Method 1 2 3 4
-GCN [33] 0.20 0.19 0.13 0.13
-Tree-LSTM [38] 1.81 2.44 1.79 1.89
-DEFT 0.76 0.70 0.91 0.93
-This section evaluates the computational efficiency of DEFT when compared to baseline methods.
-It also examines how employing the analytical gradient derived from Theorem 1 to compute (10) and the parallel programming approach detailed in Section IV-F, impacts overall computational time.
-All experiments in this section were conducted in Python on an Ubuntu 20.04 equipped with an AMD Ryzen PRO 5995WX CPU, 256GB of RAM, and 128 cores.
-V-D1 Computational Speed Comparison
-We compare computational speed for one step inference.
-Table III indicates that GCN has the fastest speed.
-DEFT is the second fastest and is able to generate single step predictions at a 100Hz frequency.
-V-D2 Computational Speed with Analytical Gradient
-We compare the computational speed of using analytical gradients when compared to using numerical gradients.
-As summarized in Table IV, computing (10) with the analytical gradient is approximately twice as fast as using numerical gradients.
-TABLE IV: Compuational Time for (10) with Dataset 1
-Time ($10^{-2}$ s)
-Method 1 2 3 4
-Analytical Gradient 0.34 0.27 0.41 0.46
-Numerical Gradient 0.74 0.65 0.84 0.91
 Figure 12: Visualization of planning for BDLO 1 for two manipulation tasks, using DEFT, a DEFT ablation that leaves out the constraint described in Theorem 5, and Tree-LSTM.
@@
 The DEFT model enables the planning algorithm to successfully complete the task, whereas the ablation approach of DEFT and the Tree-LSTM model both fail to finish tasks.
-V-D3 Computational Speed with Parallel Programming
-We compare the computational performance of using parallel programming approach described in Section IV-F to a naive (sequential) implementation, in which each branch is simulated one after the other.
-In this experiment, we use the material properties of BDLO1 and gradually increase the number of children branches from 1 to 8.
-As shown in Figure 9, the naive implementation’s computation time grows more rapidly than that of the parallel approach.
-However, a certain increase in computational time is still expected with parallelization.
-As illustrated in Figure 3, constraints at each junction between the parent branch and its child branches prevent a fully batch-based operation, thereby resulting in an increase in the computational time.
-TABLE V: Ablation Study with Dataset 1
-Modeling Accuracy (RSME, $10^{-2}$ m)$\downarrow$
-Dataset 1 Dataset 2
-Method 1 2 3 4 1 2
-W/O Constraints.2, 3 and 4 $7.66\times 10^{8}$ $6.10\times 10^{8}$ $6.56\times 10^{6}$ $1.77\times 10^{8}$ $3.81\times 10^{7}$ $1.23\times 10^{8}$
-W/O Constraints.3 and 4 $1.36\times 10^{2}$ $1.81\times 10^{2}$ $1.51\times 10^{2}$ $7.64\times 10^{1}$ $8.72\times 10^{1}$ $3.38\times 10^{2}$
-W/O Constraint.4 3.95 3.49 2.03 2.15 4.01 4.88
-W/O Learning + System ID 2.46 3.53 1.85 1.77 2.03 2.21
-W/O System ID 2.24 3.00 1.67 1.50 1.66 1.95
-W/O Residual Learning 1.93 3.19 1.70 1.62 1.86 1.89
-Full Model 1.87 2.82 1.51 1.41 1.49 1.80
-Figure 9: Computational speed comparison: parallel programming versus a naive (sequential) implementation.
-V-E Ablation Study
-The results presented in Table V demonstrate the contribution of each DEFT component.
-Notably, excluding Theorems 2, 3, and 4 causes a significant increase in prediction loss.
-This outcome is primarily driven by the absence of Theorem 2, which leads to simulation instability.
-Although residual learning is used to mitigate errors, directly modeling stiff behaviors, such as the inextensibility of DLOs, remains challenging, making the learning process highly sensitive to inputs and hampering effective gradient propagation.
-A key challenge in using neural networks to learn position-based constraints is that constraints such as inextensibility confine the system to a low-dimensional manifold within a high-dimensional space.
-Without these constraints, it becomes difficult for the network alone to keep the system on that manifold.
-Over multiple open-loop predictions, even minor errors can accumulate and push the system state off the manifold, causing reduce of residual learning’s effect.
-Similarly, removing Theorem 3 allows the children branch to detach from parent branch, and therefore fall freely under gravity, leading to a substantial increase in loss.
-Moreover, excluding Theorem 4 reduces accuracy, demonstrating the importance of introducing rigidity at junctions to allow dynamic propagation across branches.
-Despite these challenges, auto-tuning the material properties provides the most substantial improvement over the baseline, illustrating DEFT’s ability to leverage real-world data for parameter identification.
-Residual learning also yields a non-trivial performance gain.
-By contrast, the model that do not learn from real-world data underperforms those that do, illustrating the importance of differentiability in this framework.
-V-F Planning Results
-This section demonstrates how these BDLO models can be used for planning tasks, including shape matching and thread insertion.
-We compare three approaches: DEFT, which is our proposed method; DEFT without the constraint described in Theorem 4, an ablation approach that removes rigidity at junctions and thus limits dynamic propagation across branches; Tree-LSTM, the baseline model with performance closest to DEFT.
-In these experiments, we employ ARMOUR [39], a receding-horizon trajectory planner and tracking controller, to manipulate BDLOs.
-ARMOUR relies on each model’s predicted BDLO states to guide the BDLO to a desired final configuration.
-Further details on integrating ARMOUR with BDLO simulators are provided in Appendix -E.
-The results are summarized in Table VI, and Figure 8 shows one of the real-world results.
-A video of these experiments can be found in the supplementary material.
-TABLE VI: Real-World Trajectory Planning Results
-Successful Rate
-Shape Matching Thread Insertion
-Tree-LSTM [38] 2/40 0/35
-DEFT w/o Theorem 4 27/40 16/35
-DEFT 35/40 29/35
-V-F1 Shape Matching
-In this experiment a goal configuration for the wire is specified and a random initial condition for the wire is chosen.
-A trial is considered successful if the Euclidean distance between each ground-truth vertex and its corresponding target vertex after planning is less than 0.05 m.
-If the task is not completed within 30 seconds, the trial is marked as a failure.
-As summarized in Table  VI (left column), DEFT achieves a higher success rate than the other two models.
-Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
-V-F2 Thread Insertion
-In this experiment, the planning algorithm must insert the end of the BDLO into a hole of radius $2.5$cm which is placed in arbitrary location.
-The process consists of two stages—shape matching and thread insertion—which together require two goal configurations for the wire.
-Unlike the previous experiment, in which the robot grasps the branch end, the robot’s end effector clamps the third vertex from the branch’s insertion point.
-The trial is deemed successful if the BDLO branch successfully passes through the torus while also matching the user-specified wire shape successfully.
-As shown in Table  VI (right column), DEFT again achieves a higher success rate than the other two models.
-Note that these experiments were all conducted in the real-world which further demonstrates the practical applicability of DEFT.
-VI Limitations
-While DEFT shows promising results in modeling and control of BDLOs, several limitations remain that could be addressed in future work.
-First, DEFT currently lacks a contact model for accurately predicting interactions between deformable components and their surroundings.
-Incorporating a contact modeling framework, such as that proposed in [40], would enable more realistic simulations and more robust planning in environments where collisions and frictional forces play a significant role.
-Second, for the insertion tasks considered in this work, we assume that the location and geometry of the target hole are known in advance.
-In unstructured or partially observed environments, this assumption may not hold.
-Incorporating 3D scene understanding methods, such as Gaussian Splatting [41], could help autonomously detect or refine the location of target features and thereby increasing the system’s applicability to real-world scenarios.
-Lastly, this work does not account for the orientation of the gripper during BDLO manipulation, which can significantly affect deformation behavior.
-As discussed in [16], gripper orientation can influence the kinematics of deformable objects.
-Future extensions of DEFT should therefore incorporate a more comprehensive treatment of gripper orientation to improve accuracy and expand the range of manipulable tasks.
-VII Conclusion
-This paper introduces DEFT, a method that embeds residual learning within a novel differentiable branched Deformable Linear Object (BDLO) simulator.
-DEFT enables accurate modeling and real-time prediction of a BDLO’s dynamic behavior over long time horizons under dynamic manipulation.
-To demonstrate DEFT’s efficacy, we conduct a comprehensive set of experiments evaluating its accuracy, computational speed, impacts of its constribution and utility for manipulation tasks.
-Compared to the state of the art, DEFT achieves higher accuracy while remaining sample-efficient and without compromising computational speed.
-This paper also illustrates how to integrate DEFT with a planning and control framework for 3D shape matching and thread insertion of a BDLO in real-world scenarios.
-DEFT successfully completes more complex, real-world manipulation tasks than other tested baselines.
-Acknowledgement
-The authors would like to gratefully thank the support by Ford Motor Company.
+Appendix D ARMOUR
+To address the shape matching task, we employ ARMOUR [33], an optimization-based framework for motion planning and control.
+The objective is to guide a robot arm so that it manipulates a BDLO from an initial configuration to a specified target configuration.
+ARMOUR accomplishes this through a receding-horizon approach: at each iteration, it solves an optimization problem to determine the robot’s trajectory.
+ARMOUR parameterizes the robot’s joint trajectories as polynomials, with polynomial coefficients serving as the decision variables.
+One end of the rope is rigidly attached to the robot’s end effector, whose pose is determined by forward kinematics.
+By substituting the end effector’s final pose into DEFT, we obtain a prediction of the rope’s configuration at the conclusion of the motion.
+The optimization’s cost function then minimizes the Euclidean distance between this predicted BDLO configuration from DEFT and the target configuration.
+To ensure the resultant motion is physically feasible, ARMOUR incorporates constraints on the robot’s joint positions, velocities, and torques.
+Once the optimization is solved, the resulting trajectory is tracked by ARMOUR’s controller.
+This planning–execution process is repeated until the rope achieves the desired configuration or a maximum iteration threshold is exceeded (in which case the task is deemed a failure).
+Additional details on ARMOUR’s trajectory parameterization and its closed-loop controller can be found in [33, Section IX].